### v1

# Fast Revenue Maximization

Abstract
We study a data-driven problem pricing problem in which a seller offers a price for a single item based on demand observed at a small finite number of historical prices. Our goal is to derive precise evaluation procedures of the value of the historical information gathered by the seller, along with prescriptions for more efficient price experimentation. Our main methodological result is an exact characterization of the maximin ratio (defined as the worst-case revenue garnered by a seller who only relies on past data divided by the optimal revenue achievable with full knowledge of the distribution of values). This result allows to measure the value of any historical data consisting of prices and corresponding conversion rates. We leverage this central reduction to provide new insights about price experimentation. Motivated by practical constraints that impede the seller from changing prices abruptly, we first illustrate our framework by evaluating the value of local information and show that the mere sign of the gradient may sometimes provide significant information to the seller. We then showcase how our framework can be used to run efficient price experiments. On the one hand, we develop a method to select the next price experiment that the seller should use to maximize the information obtained. On the other hand, we demonstrate that our result allows to considerably reduce the price experimentation needed to reach preset revenue guarantees through dynamic pricing algorithms.
Keywords. pricing, data-driven algorithms, robust pricing, value of data, distributionally robust optimization, conversion rate, limited information.
1 Introduction and Problem Formulation
In this work, we investigate the monopoly pricing problem in which a decision-maker needs to select a price to sell a single good to a consumer. This operational problem crystallizes the trade-off between volume of sales and profit margin as setting a price too low hurts the margin but increases the demand for the product whereas setting it too high has the reverse effect.
An important practical challenge in pricing is that the distribution of willingness-to-pay of consumers, i.e., the largest price at which a consumer is willing to purchase the good, is unknown to the decision-maker. Therefore, decision-makers need to rely on historical information which commonly includes past prices offered and consumers’ purchase decisions at the offered price. Our goal in this work is to provide a principled framework for two complementary tasks: i.) quantify the value of classical data structures in the form of prices and associated conversion rates, and ii.) perform efficient price experimentation.
1.1 Problem formulation
We consider the problem of a seller trying to sell one indivisible good to one buyer. We assume that the buyer’s value $v$ is drawn from some unknown distribution $F$ with support included in $[\underline{v},\overline{v}]$, where $0\leq\underline{v}<\overline{v}<\infty$. The seller knows that the distribution of values belongs to some non-parametric class of distributions $\pazocal{C}$. The seller also has information about conversion rates observed at historical prices offered. Formally, let $\mathbf{p}=(p_{i})_{i\in\{0,\ldots N+1\}}$ be a non-decreasing sequence of prices such that $p_{0}=\underline{v}$ and $p_{N+1}=\overline{v}$ and let $\mathbf{q}=(q_{i})_{i\in\{0,\ldots N+1\}}$ be a sequence of historical conversion rates observed such that $q_{0}=1$ and $q_{N+1}=0$. The decision-maker information set is defined as
$\mathscr{I}_{\mathbf{p},\mathbf{q}}=\{(p_{i},q_{i});\quad i=0\cdots N+1\}.$
The problem we are interested in is the following: how can the seller leverage the information contained in $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ to maximize her revenue, and how should one guide price experimentation.
Pricing mechanisms and performance. The decision-maker chooses a (potentially) randomized pricing mechanism characterized by the cdf of prices the seller posts. We let P={Ψ in $\pazocal{P}=\{\Psi\mbox{ in }\pazocal{D}_{[\underline{v},\overline{v}]}\}$ be the set of randomized mechanisms, where $\pazocal{D}_{[\underline{v},\overline{v}]}$ is the set of probability measures on $[\underline{v},\overline{v}]$.Then, for every $\Psi\in\pazocal{P}$ and any distribution of consumer value $F$, the expected revenue generated from the randomized mechanism is defined as,
[…]
where we let $\mathrm{Rev}\left(p|\overline{F}\right):=p\cdot\overline{F}(p)$. The seller aims to select a mechanism whose revenue is as close as possible to the maximal revenue collected with knowledge of the distribution of values $F$. This revenue is generated by a posted price mechanism (Riley and Zeckhauser, 1983) and we define it as $\mbox{opt}(F):=\sup_{p\in[\underline{v},\overline{v}]}\mathrm{Rev}\left(p|\overline{F}\right).$ We also define the set of optimal prices as $\mathscr{O}(F)=\arg\max_{p\in[\underline{v},\overline{v}]}\mathrm{Rev}\left(p|\overline{F}\right)$.
In what follows, we evaluate the performance of a mechanism $\Psi$ against a value distribution $F$ through the ratio which compares the revenue generated by the mechanism and the optimal revenue and is defined as,
[…]
Nature’s problem. We evaluate the performance of a given pricing mechanism $\Psi$ by considering its worst-case ratio over a class of admissible value distributions denoted as $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ and defined as,
$\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})=\left\{F\in\pazocal{C}\text{ s.t. }\overline{F}(p_{i})=q_{i}\text{ for all }(p_{i},q_{i})\in\mathscr{I}_{\mathbf{p},\mathbf{q}}\right\}.$
Hence, the set $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ contains all value distributions in $\pazocal{C}$, which are consistent with the information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ observed by the decision-maker. In turn, we define Nature’s problem as,
$\inf_{F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}\mathrm{Ratio}\left(\Psi,F\right).$ (1.2)
Given a pricing mechanism $\Psi$, the value of (1.2) is the worst-case performance of the mechanism over any distribution agreeing with $\mathscr{I}_{\mathbf{p},\mathbf{q}}$. A higher ratio indicates that the information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ is highly informative in terms of revenue maximization.
Finally, we note that (1.2) quantifies the performance of a fixed policy. In Section 3, we study the value of certain information sets by computing the best achievable worst-case ratio for any pricing mechanism, which is formally defined as,
$\mathfrak{R}\left(\pazocal{P},\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\right):=\sup_{\Psi\in\pazocal{P}}\inf_{F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}\mathrm{Ratio}\left(\Psi,F\right).$ (Maximin)
In this work, we derive results on the performance of pricing mechanisms for two classes of distributions. The class of General distributions denoted as $\pazocal{G}$ that contains all distributions supported on $[\underline{v},\overline{v}]$. We also consider the subset of $\pazocal{G}$ denoted as $\pazocal{F}$ and which denotes the set of regular distributions, i.e., distributions that admit a density, except potentially at the boundaries of their support and that have a non-decreasing virtual value $v-\overline{F}(v)/f(v)$. We note that the regular class of distributions contains a wide variety of distributions; see, e.g., Ewerhart (2013) for a summary of widely used regular distributions. This class is central to the pricing and mechanism design literature and can be alternatively described as the class of distributions that induce a concave revenue function in the quantity space.
1.2 Contribution
Performance evaluation for any pricing mechanism. To evaluate the performance of different pricing mechanisms one needs to solve the optimization problem defined in (1.2). Doing so is a priori daunting as (1.2) is an constrained infinite-dimensional optimization problem over the space $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ which is a non-parametric subset of value distributions. The constraints arises in the form of shape constraints (e.g. when $\pazocal{C}=\pazocal{F}$ and distributions have to be regular) or constraints on quantiles of the distribution. Our first main contribution consists in reducing the problem (1.2) to one that is amenable to optimization. Formally, we show the following.
Theorem 1 (Reduction for Policy Evaluation).
Assume $\pazocal{C}$ is equal to $\pazocal{G}$ or $\pazocal{F}$. Fix $N\geq 1$ and let $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ be an information set including $N$ historical prices and conversions rates. Assume $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-empty then, for any mechanism $\Psi$ in $\pazocal{P}$,
$\inf_{F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}\mathrm{Ratio}\left(\Psi,F\right)=\inf_{{r^{*}}\in\mathscr{S}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}\mathrm{Ratio}\left(\Psi,F_{\pazocal{C}}(\cdot|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}})\right),$
where $F_{\pazocal{C}}(\cdot|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is a distribution defined in (2.10) and $\mathscr{S}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is a subset of $[\underline{v},\overline{v}]$ described in Proposition 2.
Theorem 1 reduces the initial infinite dimensional optimization problem in (1.2) to a one-dimensional one for any pricing mechanism. It can be used to evaluate and compare pricing-mechanisms under different assumptions about the information the decision-maker has at hand.
We present the proof of Theorem 1 in Section 2. This result significantly generalizes (Allouah et al., 2023, Theorem 1) across several dimensions. First, our reduction also applies when $\pazocal{C}$ is the set of general distributions and therefore allows to investigate settings where the decision-maker does not impose shape constraints. More importantly, our reduction holds for information sets with arbitrary numbers of historical prices $N$ while the analysis in Allouah et al. (2023) only holds for $N=1$. This considerably extends the possible information sets $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ that we can explore and, as we will see, allows us to derive novel insights on price experimentation. We also note that while this work emphasizes on the ratio objective, our proof readily extends to the absolute regret.
New insights on price experimentation. Theorem 1 provides a principled approach to: $i)$ evaluate historical demand observations, $ii)$ prescribe efficient methods to run price experiments.
First, we illustrate in Section 3.1 how our method can be used to shed light on the value of certain type of price experiments. We consider the practical setting where the seller cannot substantially vary the price offered and therefore observed demand realizations at two neighboring prices. We show that our method provides a rigorous way to assess the value of these local experiments and enable us to precisely quantify the value of the gradient as a function of its magnitude. Quite notably, we show that when the demand is regular, the sign of the gradient may already provide considerable information to the seller.
We then use our framework as a way to develop or enrich algorithms for price experimentation. In Section 3.2, we consider a setting in which the seller observed the demand at a single historical point and then may choose a new price to experiment before committing to a revenue maximizing price. We show that by carefully selecting the second price to experiment with, the seller can significantly increase the worst-case ratio, by $10$ to $20\%$ across all possible demand observed at the first price. In addition, in Section 3.3 we demonstrate that our framework can be used to improve dynamic pricing algorithms. By better quantifying the value of data, one can reduce the extent of price experimentation needed to reach a certain revenue guarantee. We consider the case in which the value distribution is regular (and hence the revenue curve is unimodal), and compare our stopping criterion with the commonly used one when experimenting with the Ternary Search algorithm. Our criterion allows to divide by $2$ to $4$ the number of prices offered to ensure a ratio larger than $99\%$. Hence, our procedure can significantly reduce the cost related to price experimentation.
1.3 Literature Review
Pricing when facing an unknown value distribution has been first considered in a distributionally robust framework with support information (Bergemann and Schlag, 2008; Eren and Maglaras, 2010) and then analyzed with moment information (Chen et al., 2022). Variants of this problem were analyzed when customer commits to a price scheme on a selling horizon (Caldentey et al., 2016) or when the decision-maker chooses among a broader set of mechanisms (see Carroll (2019); Chen et al. (2023); Wang et al. (2024) and references therein).
The pricing problem has then been studied in the offline data-driven setting where the decision-maker has past data about the value distribution and makes a single pricing decision. In the case where past data takes the form of i.i.d. values sampled from the value distribution, Huang et al. (2015) quantifies the number of samples required to be within a factor $(1-\epsilon)$ from the optimal revenue achievable with full knowledge of the distribution. Fu et al. (2015); Babaioff et al. (2018); Daskalakis and Zampetakis (2020); Allouah et al. (2022) provide performance guarantees when the seller has access to limited number of samples.
The data structure we consider in this work is more closely related to the one studied in the dynamic pricing literature where the decision-maker observes noisy demand observation at an offered price. Kleinberg and Leighton (2003); Besbes and Zeevi (2009); Broder and Rusmevichientong (2012); Keskin and Zeevi (2014); Singla and Wang (2023) study the performance of dynamic pricing policies trading-off between exploration and exploitation in parametric and non-parametric settings. Our work differs from this previous literature along various dimensions: we measure the performance of pricing algorithms with respect to the a single-shot pricing decision (as in Leme et al. (2023b)) as opposed to the commonly studied cumulative regret, hence there is no tradeoff between exploration and exploitation in our work. Furthermore, we derive an exact characterization of the worst-case performance of algorithms that allows to exactly quantify the value of past data. All the previous work focuses on rate-optimal guarantees. Finally, our work assumes that the decision-maker has access to non-noisy demand observation.
Our work is more closely related to Allouah et al. (2023), which provides a tractable procedure to arbitrarily approximate the maximin ratio when $N=1$. Our work extends their results to settings where the decision-maker has access to demand information across several prices. This allows us to derive new insights and prescriptions regarding efficient price experimentation.
Finally, we note that our work broadly relates to Meister and Nietert (2021); Leme et al. (2023a); Okoroafor et al. (2023) which study the sample complexity required to estimate the cumulative distribution functions with binary feedback and to Stoye (2009); Besbes and Mouchtaki (2023); Besbes et al. (2023) which provide tight characterizations of the value of data for other problems.
2 Proof of Theorem 1
In this section we prove our main methodological result. Assume $\pazocal{C}$ is equal to $\pazocal{G}$ or $\pazocal{F}$. Fix $N\geq 1$ and let $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ be an information set including $N$ historical prices and conversions rates. Fix a pricing mechanism $\Psi$ in $\pazocal{P}$. We first, note that one can prove the result by considering the more general notion of $\lambda$-regret defined for every $\lambda\in(0,1]$ as,
[…]
Taking $\lambda=1$ recovers the absolute regret metric. Furthermore, by using a classical epigraph reformulation (described in Appendix A) of the worst-case ratio one can reduce its analysis to the one of the $\lambda$-regret (see Proposition 1 in Anunrojwong et al. (2023)). In what follows, we prove that the reduction stated in Theorem 1 holds for the $\lambda$-regret.
The proof generalizes the reasoning developed in (Allouah et al., 2023). In Section 2.1 we construct candidate worst-case distributions for problem (1.2). In Section 2.2, we reduce the infinite-dimensional optimization problem to an optimization problem over two parameters, the optimal price of the worst-case distribution and the conversion rate associated to the optimal price. In Section 2.3 we further simplify the optimization problem by fully characterizing the optimal conversion rate for any fixed optimal price and therefore end up with a one-dimensional optimization problem.
Notations. We define for every $F$ in $\pazocal{D}_{[\underline{v},\overline{v}]}$, the inverse $F^{-1}(q):=\inf\{v\mbox{ in }[\underline{v},\overline{v}]\mbox{ s.t. }F(v)\geq q\}$ for all $q$ in $[0,1]$. For a distribution with positive density function $f$ on its support $[a,b]$, where $0\leq a<\infty$ and $a\leq b\leq\infty$, we denote the virtual value function for $v\in[a,b]$ by $\phi_{F}(v):=v-\frac{\overline{F}(v)}{f(v)}.$ For any $v\geq 0$, we define $\Gamma(v):=\frac{1}{1+v}$ and $\Gamma^{-1}(q):=\frac{1}{q}-1$ denotes the inverse of $\Gamma$.
2.1 Candidate worst-case distributions
We first introduce an important auxiliary distribution which is the central building block of our construction of the candidate worst-case distribution. For any pair of values $(s,s^{\prime})$ such that $0\leq s\leq s^{\prime}$ and $01\geq q_{s}\geq q_{s^{\prime}}>$ 0, we define on $[0,\infty)$ the following ccdfs.
$\overline{G}_{\pazocal{G}}(v|(s,q_{s}),(s^{\prime},q_{s^{\prime}}))$ $=\mathbbm{1}\{v\in[\underline{v},\overline{v}]\}\cdot q_{s^{\prime}}$ (2.1)
$\overline{G}_{\pazocal{F},}(v|(s,q_{s}),(s^{\prime},q_{s^{\prime}}))$ $=\mathbbm{1}\{v\in[\underline{v},\overline{v}]\}\cdot\Gamma\left(\Gamma^{-1}\left(q_{s}\right)+\frac{\Gamma^{-1}\left(q_{s^{\prime}}\right)-\Gamma^{-1}\left(q_{s}\right)}{s^{\prime}-s}(v-s)\right)$ (2.2)
These distributions are motivated by the fact that they allow us to derive the following bounds on distributions in $\pazocal{C}$ by using the so-called single crossing property (Barlow and Proschan, 1975).
Lemma 1 (Local Bounds).
Fix $F$ in $\pazocal{C}$ and values $(s,s^{\prime})$ such that $0\leq s\leq s^{\prime}$ and $0<q_{s^{\prime}}=\overline{F}(s^{\prime})\leq q_{s}=\overline{F}(s)$. We have that,
$\overline{F}(v)$ $\geq\overline{G}_{\pazocal{C}}(v|(s,q_{s}),(s^{\prime},q_{s^{\prime}}))\qquad\mbox{ if }\>v\mbox{ in }[s,s^{\prime}]$
$\overline{F}(v)$ $\leq\overline{G}_{\pazocal{C}}(v|(s,q_{s}),(s^{\prime},q_{s^{\prime}}))\qquad\mbox{ if }\>v>s^{\prime}\text{ or }v<s.$
Lemma 1, proved in Section 5, shows that $\overline{G}_{\pazocal{C}}$ provides a way to upper and lower bound all function in the class $\pazocal{C}$ sharing the same conversion rates at $s$ and $s^{\prime}$. Given that every distribution in $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ have the same conversion rate at prices $\mathbf{p}$, we next use Lemma 1 to construct candidate worst-case distributions. In particular, given a feasible information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}=\{(p_{i},q_{i});\quad i=0\cdots N+1\}$ we define for the sake of simple notations, $G_{\pazocal{C},i}(v)={G}_{\pazocal{C}}(v|(p_{i},q_{i}),(p_{i+1},q_{i+1}))$ for every $i\in 0\cdots N$. For both the regular and general classes of distributions, we define on $[\underline{v},\overline{v})$:
$\overline{L}_{\pazocal{C}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}})=\overline{G}_{\pazocal{C},i}(v),\quad\mbox{if }v\in[p_{i},p_{i+1}),\mbox{ for }i=0\cdots N.$ (2.3)
Furthermore, we let
$\overline{U}_{\pazocal{G}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ $=\overline{G}_{\pazocal{G},i-1}(v)=q_{i}\quad\mbox{if }v\mbox{ in}[p_{i},p_{i+1}),\mbox{for }i=0\cdots N.$ (2.4)
$\overline{U}_{\pazocal{F}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ =$=\begin{cases}\min\{1,\overline{G}_{\pazocal{F},1}(v)\}&\quad\mbox{if }v\mbox{ in }[p_{0},p_{1})\\ \min\{\overline{G}_{\pazocal{F},{i-1}}(v),\overline{G}_{\pazocal{F},{i+1}}(v)\}&\quad\mbox{if }v\mbox{ in }[p_{i},p_{i+1}),\mbox{for }i=1\cdots N-2\\ \overline{G}_{\pazocal{F},{i-1}}(v)&\quad\mbox{if }v\mbox{ in }[p_{i},p_{i+1}),\mbox{ for }i\in\{N-1,N\}.\end{cases}$ (2.5)
Intuitively, Lemma 1 implies that the ccdf $\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is a lower bound of the ccdf of any distribution in $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Similarly, the ccdf $\overline{U}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ provides an upper bound. We illustrate this in Figure 1.
a) Bounds on the CCDF values.
b) Bounds on the revenue values.
Figure 1: Illustration of the implications of the bounds obtained by the single crossing property.
2.2 Reduction to 2D-problem
In this section, we prove our first reduction and develop a reduction of nature’s problem to a two-dimensional optimization problem parameterized by the set of possible values that the optimal price and quantile can take. We first define the set of feasible values given the information at hand. To that end, let ${\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ denote the set of feasible optimal pairs, i.e.,
${\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}}):=\left\{({r^{*}},{q^{*}})\mbox{ in }\mathbb{R}_{+}\times[0,1]:\mbox{there exists }\>F\mbox{ in }\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\mbox{ s.t. }{r^{*}}\in\mathscr{O}(F)\mbox{ and }\overline{F}({r^{*}})={q^{*}}\right\}.$
Given such a definition, we have for every $\lambda\in[0,1]$,
$\sup_{F\mbox{ in }\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}R_{\lambda}\left(\Psi,F\right)=\sup_{({r^{*}},{q^{*}})\in{\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}\left\{\lambda\cdot{r^{*}}\cdot{q^{*}}-\inf_{\begin{subarray}{c}F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}):\\ {r^{*}}\in\mathscr{O}(F),\overline{F}({r^{*}})={q^{*}}\end{subarray}}\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|\overline{F}\right)]\right\}$ (2.6)
The above reformulation allows to “decouple” the problem as the first term is fully controlled and second term representing the expected revenue of the pricing mechanism can be optimized independently. Our next structural result solves exactly the inner minimization problem.
Proposition 1 (Worst-case Revenue).
Assume $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is not empty. For every $({r^{*}},{q^{*}})\in{\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ and every pricing mechanism $\Psi$, we have that,
$\inf_{\begin{subarray}{c}F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}):\\ {r^{*}}\in\mathscr{O}(F),\overline{F}({r^{*}})={q^{*}}\end{subarray}}\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|\overline{F}\right)]=\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))\right)].$
We note that Proposition 1 and (2.6) implies that the problem of nature can be written as
$\sup_{F\mbox{ in }\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}R_{\lambda}\left(\Psi,F\right)=\sup_{({r^{*}},{q^{*}})\in{\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}\mathbb{E}_{p\sim\Psi}\left[\lambda\cdot{r^{*}}\cdot{q^{*}}-\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))\right)\right]$ (2.7)
In other words, given ${r^{*}}\in\mathscr{O}(F)$ and ${q^{*}}$ s.t. ${q^{*}}=\overline{F}({r^{*}})$, one can pin down the worst-case distribution and associated revenue curve.
In the rest of this section, we prove Proposition 1. Our first lemma provides a lower bound on the revenue associated with any value distribution $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$.
Lemma 2.
Consider an information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}$, then for any pricing mechanism $\Psi\in\pazocal{P}$ and any value distribution $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$, we have that, $\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|\overline{F}\right)]\geq\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)]$, where $\overline{L}_{\pazocal{C}}$ is defined in (2.3).
Proof of Lemma 2.
Fix $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ and consider $v\in[\underline{v},\overline{v}]$. We first show that, $\overline{F}(v)\geq\overline{L}_{\pazocal{C}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}}).$ Consider $i\in\{0,\ldots,N\}$ such that, $v\in[p_{i},p_{i+1}]$. We remark that $\overline{F}(p_{i})=q_{i}$ and $\overline{F}_{0}(p_{i+1})=q_{i+1}$ because $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Therefore Lemma 1 implies that,
$\overline{F}(v)\geq\overline{G}_{\pazocal{C}}(v|(p_{i},q_{i}),(p_{i+1},q_{i+1}))=\overline{L}_{\pazocal{C}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}}),$
where the last equality holds by definition of $\overline{L}_{\pazocal{C}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Hence, $\overline{F}(v)\geq\overline{L}_{\pazocal{C}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Furthermore, for any pricing mechanism $\Psi\in\pazocal{P}$, we have that,
$\mathbb{E}_{\Psi}[Rev\left(p|\overline{F}\right)]=\int_{0}^{\infty}p\cdot\overline{F}(p)\,d\Psi(p)\stackrel{{\scriptstyle(a)}}{{\geq}}\int_{0}^{\infty}p\cdot\overline{L}_{\pazocal{C}}(p|\mathscr{I}_{\mathbf{p},\mathbf{q}})\,d\Psi(p)=\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)],$
where $(a)$ holds because $\overline{F}(v)\geq\overline{L}_{\pazocal{C}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}})$. This concludes the proof. ∎
Next, we show that our candidate worst-case distribution belongs to $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$.
Lemma 3.
If $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-empty, then $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$.
Proof of Lemma 3.
First, one can verify that for every $i\in\{0,\ldots,N+1\}$, $\overline{L}_{\pazocal{C}}(p_{i}|\mathscr{I}_{\mathbf{p},\mathbf{q}})=q_{i}.$ Therefore, we need to show that, $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\in\pazocal{C}.$
When $\pazocal{C}=\pazocal{G}$, we note that $L_{\pazocal{G}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-decreasing because $(q_{i})_{i\leq N+1}$ is non-decreasing. Indeed, $(q_{i})_{i\leq N+1}$ because the set $\pazocal{G}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-empty. Hence, $L_{\pazocal{G}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\in\pazocal{G}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$.
When $\pazocal{C}=\pazocal{F}$, we show in Proposition 3 (stated and proved in Section 5) that if $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-empty then $\left(\frac{\Gamma^{-1}\left(q_{i+1}\right)-\Gamma^{-1}\left(q_{i}\right)}{p_{i+1}-p_{i}}\right)_{0\leq i\leq N-1}$ is non-decreasing. We next show that this implies that $\overline{L}_{\pazocal{F}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is regular. Lemma 6 implies that the associated virtual value function satisfies
$\phi_{L_{\pazocal{F}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})}(v)={p_{i}}-({p_{i+1}}-{p_{i}})\frac{1+{\Gamma^{-1}\left(q_{i}\right)}}{{\Gamma^{-1}\left(q_{i+1}\right)}-{\Gamma^{-1}\left(q_{i}\right)}},\quad\mbox{if }v\in[p_{i},p_{i+1}),\mbox{ for }i=0\cdots N$
Hence, the virtual value function is piece-wise constant. Now we need to show that $\phi_{L}(v)$ is non-decreasing. Fix $11\leq i\leq N-$ 1. we evaluate the difference between the two consecutive constant values that the virtual value function is and note that it is equal to,
$\left(1+{\Gamma^{-1}\left(q_{i}\right)}\right)\left[\frac{{p_{i}}-{p_{i-1}}}{{\Gamma^{-1}\left(q_{i}\right)}-{\Gamma^{-1}\left(q_{i-1}\right)}}-\frac{{p_{i+1}}-{p_{i}}}{{\Gamma^{-1}\left(q_{i+1}\right)}-{\Gamma^{-1}\left(q_{i}\right)}}\right]\stackrel{{\scriptstyle(a)}}{{\geq}}0,$
where inequality $(a)$ holds because the sequence $\left(\frac{\Gamma^{-1}\left(q_{i+1}\right)-\Gamma^{-1}\left(q_{i}\right)}{p_{i+1}-p_{i}}\right)_{0\leq i\leq N-1}$ is non-decreasing.
Hence, the virtual value function of $\overline{L}_{\pazocal{F}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-decreasing and $L_{\pazocal{F}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\in\pazocal{F}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. ∎
The final structural result needed in the proof of Proposition 1 characterizes the optimal revenue (and associated conversion rate) of $\overline{L}_{\pazocal{C}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}})$.
Lemma 4.
If $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-empty, then $\sup_{p\in[\underline{v},\overline{v}]}\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)=\max_{i\in\{0,\ldots,N+1\}}p_{i}\cdot q_{i}.$
Proof of Lemma 4.
We first note that, Lemma 3 implies that $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ belongs to $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Therefore, for every $i\in\{1,\ldots,N+1\}$, we have $\mathrm{Rev}\left(p_{i}|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)=p_{i}\cdot\overline{L}_{\pazocal{C}}(p_{i}|\mathscr{I}_{\mathbf{p},\mathbf{q}})=p_{i}\cdot q_{i}$. Hence, $\sup_{p\in[\underline{v},\overline{v}]}\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)\geq\max_{i\in\{0,\ldots,N+1\}}p_{i}\cdot q_{i}.$
To prove the reverse inequality, it suffices to show that the revenue function $p\mapsto\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)$ is monotonic on every interval $[p_{i},p_{i+1})$ for $i\in\{0,\ldots,N\}$. By the definition in (2.3), we need to analyze the monotonicity of the function $\kappa:p\mapsto p\cdot\overline{G}_{\pazocal{C},i}(p)$.
When $\pazocal{C}=\pazocal{G}$, we have that $\overline{G}_{\pazocal{C},i}(p)$ is constant on $[p_{i},p_{i+1})$. Hence, $\kappa$ is a linear function and is monotonic.
When $\pazocal{C}=\pazocal{F}$, we define for every $q\in[\overline{G}_{\pazocal{C},i}(p_{i}),\overline{G}_{\pazocal{C},i}(p_{i+1})]$ as $\tilde{\kappa}(q)=\overline{G}_{\pazocal{C},i}^{-1}(q)\cdot q$. We note that for every $p\in[p_{i},p_{i+1}]$ we have $\kappa(p)=\tilde{\kappa}(\overline{G}_{\pazocal{C},i}(p))$ and, by composition, if $\tilde{\kappa}$ is monotonic then $\kappa$ is monotonic. To show that $\tilde{\kappa}$ is monotonic we note that it corresponds to the revenue function in the quantile space. Its derivative satisfies, $\frac{d}{dq}\tilde{\kappa}(q)=\phi_{G_{\pazocal{C},i}}(\overline{G}_{\pazocal{C},i}^{-1}(q))$, where $\phi_{F}$ is the virtual value function of a distribution $F$. To conclude the proof, we show in Lemma 6 that the virtual value function of $G_{\pazocal{C},i}$ is constant. Therefore, $\tilde{\kappa}$ is linear and monotonic. ∎
We conclude this section with the proof of Proposition 1.
Proof of Proposition 1.
Consider $({r^{*}},{q^{*}})\in{\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ and consider the extended information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+}:=\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup\{({r^{*}},{q^{*}})\}$. We note that $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$ is the subset of $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ where distributions also satisfy $\overline{F}({r^{*}})={q^{*}}$. We also let $\mathscr{A}=\{F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\text{ s.t. }{r^{*}}\in\mathscr{O}(F)\text{ and }\overline{F}({r^{*}})={q^{*}}\}.$ We first show that,
$\inf_{F\in\mathscr{A}}\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|\overline{F}\right)]\geq\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\right)].$ (2.8)
Indeed by definition, we have that $\mathscr{A}\subset\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$. Therefore, (2.8) follows from Lemma 2 applied to the information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+}$. Next, we show that
$\inf_{F\in\mathscr{A}}\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|\overline{F}\right)]\leq\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\right)].$ (2.9)
To do so, we will show that $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\in\mathscr{A}$ by proving that $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ and that $({r^{*}},{q^{*}})$ are optimal prices and associated conversion rates for $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$.
We first show that $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$ is non-empty. Note that since $({r^{*}},{q^{*}})\in{\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ there exists $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ such that $\overline{F}({r^{*}})={q^{*}}$. Such a distribution belongs to $\mathscr{A}$ and therefore belongs to $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$. Therefore, this set is not empty. It then follows from Lemma 3 that $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\subset\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$.
In addition, Lemma 4 applied to the information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+}$ implies that,
$\sup_{p\in[\underline{v},\overline{v}]}\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\right)=\max\left(\max_{i\in\{0,\ldots,N+1\}}p_{i}\cdot q_{i},{r^{*}}\cdot{q^{*}}\right).$
Furthermore, we argue that ${r^{*}}\cdot{q^{*}}\geq\max_{i\in\{0,\ldots,N+1\}}p_{i}\cdot q_{i}$. This inequality holds because $\mathscr{A}$ is non-empty and for any distribution $F\in\mathscr{A}$, we have for every $i\in\{0,\ldots,N+1\}$ that ${r^{*}}\cdot{q^{*}}={r^{*}}\cdot\overline{F}({r^{*}})\stackrel{{\scriptstyle(a)}}{{\geq}}p_{i}\cdot\overline{F}(p_{i})=p_{i}\cdot q_{i},$ where $(a)$ holds because $F\in\mathscr{A}$ and hence ${r^{*}}$ is an optimal price for this distribution. This implies that $\sup_{p\in[\underline{v},\overline{v}]}\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\right)={r^{*}}\cdot{q^{*}}$ and given that $\overline{L}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})={q^{*}}$, it concludes that ${r^{*}}$ is an optimal price for $\overline{L}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$ with associated conversion rate ${q^{*}}$.
Therefore, $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\in\mathscr{A}$ and (2.9) holds. Finally, (2.8) and (2.9) imply Proposition 1. ∎
2.3 Reduction to a 1D problem
In a second step, we further reduce the minimization problem stated in Equation (2.7) to a one-dimensional minimization problem by deriving for every value of ${r^{*}}$ the conversion-rate ${q^{*}}$ which maximizes the $\lambda$-regret. For any ${r^{*}}$, we define ${\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}}):=\{q^{*}\in[0,1]\text{ s.t. }(r^{*},q^{*})\in{\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\}.$ Furthermore, we fix $\lambda\in[\mathfrak{R}\left(\pazocal{P},\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\right),1]$ and define the mapping,
$\rho:\begin{cases}[0,1]\to\mathbb{R}_{+}\\ {q^{*}}\mapsto\mathbb{E}_{p\sim\Psi}\left[\lambda\cdot{r^{*}}\cdot{q^{*}}-\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))\right)\right].\end{cases}$
We next show that $\rho$ is monotonic.
Lemma 5.
Fix $\lambda\in[\mathfrak{R}\left(\pazocal{P},\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\right),1]$. Then for every $r^{*}$ such that, ${\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-empty, the mapping $\rho$ is non-decreasing.
Proof of Lemma 5.
We note that for every $q^{*}>0$, $\rho(q^{*})=q^{*}\cdot\frac{\rho(q^{*})}{q^{*}}$. To conclude the proof it is sufficient to show that $\frac{\rho(q^{*})}{q^{*}}$ is non-negative and non-decreasing as the product of two non-negative and non-decreasing functions is non-decreasing.
First, we remark from the epigraph formulation in Appendix A that $\frac{\rho(q^{*})}{q^{*}}$ is non-negative. To prove that $\frac{\rho(q^{*})}{q^{*}}$ is non-decreasing remark that, $\frac{\rho({q^{*}})}{{q^{*}}}=\lambda r^{*}-\mathbb{E}_{p\sim\Psi}\left[\frac{1}{q^{*}}\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))\right)\right]$. Hence, it is sufficient to show that for every $p$, the mapping $\tilde{\rho}:{q^{*}}\mapsto\frac{1}{{q^{*}}}\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))\right)$ is non-increasing.
When $\pazocal{C}=\pazocal{G}$, $\tilde{\rho}$ is a constant when $\overline{L}_{\pazocal{C}}(p|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))={q^{*}}$ and is the inverse function when $\overline{L}_{\pazocal{C}}(p|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))=q_{i}\neq{q^{*}}$. In both cases, it is non-increasing.
When $\pazocal{C}=\pazocal{F}$, let $\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+}=\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}})$ be the extended information set and reindex the elements such that $\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+}=\{(\hat{p}_{i},\hat{q}_{i});\;i=0\ldots N+2\}$. Fix $p$ and let $i$ be such that $p\in[\hat{p}_{i},\hat{p}_{i+1})$ By definition of $L_{\pazocal{C}}$ we have that
$\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\right)=p\cdot\overline{G}_{\pazocal{C}}(p|(\hat{p}_{i},\hat{q}_{i}),(\hat{p}_{i+1},\hat{q}_{i+1})).$
Hence, $\mathrm{Rev}\left(p|\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\right)$ only varies as a function of ${q^{*}}$ if ${r^{*}}\in\{\hat{p}_{i},\hat{p}_{i+1}\}$. In all other cases, the numerator in $\tilde{\rho}$ is a constant which does not depend on ${q^{*}}$ and therefore $\tilde{\rho}$ is clearly non-increasing. We next show that when either $p_{i}={r^{*}}$ or $p_{i+1}={r^{*}}$, $\tilde{\rho}$ is also non-increasing.
Case 1: $p_{i}={r^{*}}$. In this case, we remark using the alternative form of $\bar{G}_{\pazocal{C}}$ derived in the proof of Lemma 1 that,
$\tilde{\rho}({q^{*}})=\frac{1}{{q^{*}}}\cdot p\cdot{q^{*}}\cdot\Gamma\left(\Gamma^{-1}\left(\frac{q_{i+1}}{{q^{*}}}\right)\frac{p-{r^{*}}}{p_{i+1}-{r^{*}}}\right)=p\cdot\Gamma\left(\Gamma^{-1}\left(\frac{q_{i+1}}{{q^{*}}}\right)\frac{v-{r^{*}}}{p_{i+1}-{r^{*}}}\right).$
By composition, we conclude that $\tilde{\rho}$ is non-increasing.
Case 2: $p_{i+1}={r^{*}}$. Using again one of the alternative forms of $\bar{G}_{\pazocal{C}}$ derived in the proof of Lemma 1 we obtain that,
$\tilde{\rho}({q^{*}})=\frac{1}{{q^{*}}}\cdot\frac{p}{\frac{1}{q_{i}}+\left(\frac{1}{{q^{*}}}-\frac{1}{q_{i}}\right)\frac{p-p_{i}}{{r^{*}}-p_{i}}}=\frac{p}{\frac{{q^{*}}}{q_{i}}\cdot\left(1-\frac{p-p_{i}}{{r^{*}}-p_{i}}\right)+\frac{p-p_{i}}{{r^{*}}-p_{i}}}.$
By composition, we conclude again that $\tilde{\rho}$ is non-increasing. ∎
To complete the reduction to a one-dimensional problem, we remark that,
$\sup_{F\mbox{ in }\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}R_{\lambda}\left(\Psi,F\right)=\sup_{r^{*}\in[\underline{v},\overline{v}]}\sup_{q^{*}\in{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}\rho(q^{*})$ =$\stackrel{{\scriptstyle(a)}}{{=}}\sup_{r^{*}\in[\underline{v},\overline{v}]}\mathbbm{1}\{{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\neq\emptyset\}\cdot\rho(\sup{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}}))$
=$\stackrel{{\scriptstyle(b)}}{{=}}\sup_{r^{*}\in[\underline{v},\overline{v}]}\mathbbm{1}\{{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\neq\emptyset\}\cdot\rho(\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})).$
Where $(a)$ follows from Lemma 5 and $(b)$ follows from Lemma 7, stated and proved in Section 5.
For every ${r^{*}}\in[\underline{v},\overline{v}]$, we consider the function defined for every $v\in[\underline{v},\overline{v}]$ as,
$\overline{F}_{\pazocal{C}}(v|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}}):=\overline{L}_{\pazocal{C}}\left(v|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}}))\right).$ (2.10)
We note that we have established that,
$\sup_{F\mbox{ in }\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}R_{\lambda}\left(\Psi,F\right)=\sup_{{r^{*}}\in[\underline{v},\overline{v}]}\mathbbm{1}\{{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\neq\emptyset\}\cdot R_{\lambda}\left(\Psi,F_{\pazocal{C}}(\cdot|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)=\sup_{{r^{*}}\in\mathscr{S}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}R_{\lambda}\left(\Psi,F_{\pazocal{C}}(\cdot|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}})\right),$
where the last equality holds by defining $\mathscr{S}(\mathscr{I}_{\mathbf{p},\mathbf{q}})=\{{r^{*}}\in[\underline{v},\overline{v}]\text{ s.t. }{\mathscr{}B}_{{r^{*}}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\neq\emptyset\}$. This concludes our reduction to a one-dimensional optimization problem.
Evaluation of the condition: ${\mathscr{}B}_{{r^{*}}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\neq\emptyset$. When performing a line search by using the reduction we developed, one needs to assess whether ${r^{*}}\in\mathscr{S}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Our next result shows that this condition can be verified efficiently.
Proposition 2.
For every ${r^{*}}\in[\underline{v},\overline{v}]$, the ${r^{*}}\in\mathscr{S}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ if and only if,
$i)$ $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})))$ is non-empty, and $ii)$ ${r^{*}}\cdot\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})\geq\max_{i\in\{0,\ldots,N+1\}}p_{i}\cdot q_{i}$.
Proof of Proposition 2.
First, if ${\mathscr{}B}_{{r^{*}}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$, Lemma 7 implies that $\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})\in{\mathscr{}B}_{{r^{*}}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. By definition, of ${\mathscr{}B}_{{r^{*}}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ this implies that $i)$ and $ii)$ are satisfied.
Reciprocally, $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})))$ is non-empty, Lemma 3 implies that the distribution $\overline{L}_{\pazocal{C}}\left(v|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}}))\right)\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})))\subset\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Furthermore, the fact Lemma 4 along with the condition that ${r^{*}}\cdot\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})\geq\max_{i\in\{0,\ldots,N+1\}}p_{i}\cdot q_{i}$, implies that ${r^{*}}$ is an optimal price for $\overline{L}_{\pazocal{C}}\left(v|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}}))\right)$ with associated conversion-rate $\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})$. This proves that $\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})\in{\mathscr{}B}_{{r^{*}}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Hence, this set is non-empty. ∎
3 Applications to Price Experimentation
We have established a tractable procedure to evaluate the value of any historical information set that comes in the form of prices and their associated conversion rates. In the following subsections, we illustrate how Theorem 1 can be leveraged to derive insights on price experimentation. In Section 3.1, we study the value of local experimentation and consider a setting in which the decision-maker cannot substantially vary prices and therefore has only an estimate of the gradient at a given price along with the conversion rate at that price. In Section 3.2, we investigate the value of global experimentation and derive an optimal (freely selected) second price to experiment at after receiving the conversion rate from a first price. Finally, in Section 3.3, we illustrate how our exact analysis could be used to considerably reduce the number of price experiments needed to reach a certain level of optimality, when using classical algorithms.
3.1 Pricing with gradient information: the value of local experimentation
In many practical settings, the seller cannot drastically change the offered price and therefore only gather local informations. Motivated by this, we consider the setting in which the seller knows the support of the demand function $[\underline{v},\overline{v}]$ and has observed two demand points, i.e. $\mathscr{I}_{\mathbf{p},\mathbf{q}}=\{(p_{1},q_{1}),(p_{\epsilon},q_{\epsilon})\}$. The two prices are such that $p_{\epsilon}=(1+\epsilon)p_{1}$ for $\epsilon>0$. In what follows, we will parametrize this instances by the estimated gradient of the revenue function at $p_{1}$ defined as,
$\nabla_{p}Rev(p|\overline{F})(p_{1})\approx\frac{p_{\epsilon}q_{\epsilon}-p_{1}q_{1}}{p_{\epsilon}-p_{1}}:=g_{\epsilon}.$
Value of the numerical gradient. We propose to understand the value of such observations by leveraging the reduction in Theorem 1 to numerically compute the maximin ratio defined in (Maximin) when the demand function is regular (i.e., $\pazocal{C}=\pazocal{F}$) and its support is included in $[\underline{v},\overline{v}]=[1,100]$. We consider the setting where the seller experiments at an increment of $\epsilon=1\%$ of the observed price $p_{1}$ which we fix at $10$ and we consider different initial observed demand $q_{1}$ in $\{0.01,0.1,0.25,0.5,0.75,0.9,0.99\}$. For each initial $q_{1}$, we evaluate a lower bound on the maximin Ratio for a grid of 200 possible values of $g_{\epsilon}$ and report the results in Figure 2. In Appendix A, we discuss in details how we compute the lower bound on (Maximin).
a) Negative observed gradient.
b) Positive observed gradient.
Figure 2: The value of a gradient measurement for various initial settings $q_{1}$.
Figure 2 naturally demonstrates that when the absolute value of the gradient is low, the worst-case ratio is guaranteed to be high as the decision-maker is very close to the optimal price. Interestingly, relatively large values of the gradient are still very informative and allow considerable improvement compared to a single price observation. When $q_{1}=0.5$, i.e., when the first price was at the median of the value distribution, the worst-case ratio without gradient information is around $60\%$ (as we will see in Figure 3). This can increase to more that $80\%$ with a gradient with an absolute value lower or equal than $0.1$.
Value of the gradient-sign: In settings where the estimation of the absolute value of the gradient is noisy and unreliable and where the seller can only estimate the sign of the gradient, our framework allows characterizing the value of such measurement when the demand is regular as it translates to the elimination of one side of the interval of feasible optimal reserve prices. Indeed, given that the revenue function is unimodal in that case, the sign of the gradient at a price $p_{1}$ allows to assess whether the optimal price ${r^{*}}$ satisfies ${r^{*}}\geq p_{1}$ or ${r^{*}}\leq p_{1}$.
We illustrate numerically the value of a positive and negative gradient measurement in the setting where the seller knows the demand function is regular and its support is included in $[\underline{v},\overline{v}]=[1,100]$, the seller estimates the gradient sign at the observed price $p_{1}$ which we fix at $10$, and we consider different initial observed demand $q_{1}$. For each initial setting $q_{1}$, we evaluate the maximin Ratio for negative and positive observed gradient-sign as well as the value of the maximin Ratio without gradient information and report the values in Figure 3.
Figure 3: The value of the sign of gradient measurement for various initial settings $q_{1}$.
[…]
3.2 The value of one additional experiment: the value of global experimentation
We now relax the constraint that the decision-maker should explore locally and consider the setting in which the seller has observed a single demand point $(p_{1},q_{1})$ and has the opportunity to experiment at any second price $p_{2}$ by collecting the demand level at the latter price, before committing to a final pricing mechanism. We model the problem as a game between Nature and the seller, in which the seller commits to a price $p_{2}$ to experiment at, and Nature counters the seller by choosing the associated feasible quantile $q_{2}$ to yield the worst maximin ratio in the final stage. More specifically, for a fixed $p_{2}$, Nature is solving the following problem,
$\inf_{q_{2}\in[0,1]}\sup_{\Psi\in\pazocal{P}}\inf_{F\in\pazocal{C}(\{(p_{1},p_{2}),(q_{1},q_{2})\})}\mathrm{Ratio}\left(\Psi,F\right).$
We illustrate numerically the value of one additional experiment in the setting where the seller knows the demand function is regular and its support is included in $[\underline{v},\overline{v}]=[1,100]$. The seller starts with one initial observed demand at the price $p_{1}=10$. For each initial setting $q_{1}$ and candidate experiment price $p_{2}$ from a fixed grid of $[\underline{v},\overline{v}]$ of size 200, we numerically evaluate a lower-bound on the above objective using the procedure described in Appendix A. We present in Figure 4-a, the improvement in the worst-case ratio allowed with the best experiment price $p_{2}$ chosen by the decision-maker and show in Figure 4-b what this best price is.
a) Performance with one additional experiment.
b) ”Best” candidate experiment price $p_{2}$.
Figure 4: The value of one additional global experiment for various initial settings $(p_{1},q_{1})$.
Figure 4 shows that by allowing the decision-maker to select two prices (as opposed to one), one can considerably improve the performance of the pricing policy. Indeed, while a single price already guarantees a ratio of at most $30\%$ for conversion rates between $[0.01,0.99]$, the decision-maker is ensured to obtain almost $50\%$ of the revenue by experimenting correctly at a second price. Figure 4-b also illustrates how our procedure allows us to select a second price to run efficient experimentations. For instance, it provides general guidelines regarding when the price should be deflated or inflated to gather more information: in this case, when the conversion rate at $p_{1}$ is lower (respectively higher) than approximately $0.3$ the decision-maker should consider experiment more by deflating (resp. inflating) the price.
3.3 Dynamic pricing with deterministic feedback
Section 3.2 provides an initial step towards dynamic pricing, where the decision-maker explores at one more price before committing. In this section, we use our framework to provide a criterion to stop pricing experiments while guaranteeing a certain level of performance. This problem is of significant importance as price experimentation is particularly costly for firms: ending them as soon as possible can lead to substantial savings. To illustrate this, we consider the case where the seller is dynamically adjusting the posted prices and observes deterministic true demand values at the implemented prices in each round.
Consider a general dynamic pricing algorithm $\theta$ that maps the history $\pazocal{H}_{t}$ of collected prices and associated observed deterministic demands in each iteration $t$ to the next price to post i.e $p_{t}=\theta\left(\pazocal{H}_{t}\right)$. We propose to use our framework as a stopping criterion for the price experimentation phase. More specifically, we propose to augment any dynamic pricing method $\theta$ as in Algorithm 1.
Input: Acceptable sub-optimality $\epsilon$;
Initialize $\pazocal{H}_{1}$, t =1;
while $\mathfrak{R}\left(\pazocal{P},\pazocal{C}(\pazocal{H}_{t})\right)>\epsilon$ do
       Offer price $p_{t}=\theta(\pazocal{H}_{t})$ posted by algorithm $\theta$ ;
       Update $\pazocal{H}_{t+1}=\pazocal{H}_{t}\cup\{(p_{t},q_{t}=\overline{F}(p_{t}))\}$ and increment $t$;
end while
Return $p_{t^{*}},$ where $t^{*}=\arg\max_{t}\left(p_{t}\cdot q_{t}\right)$;
Algorithm 1 Revenue Maximization Algorithm $\theta$ with stopping criterion
When the underlying value distribution is regular, i.e., $\pazocal{C}=\pazocal{F}$, the revenue function is unimodal. In this case, a popular algorithm is the Ternary search algorithm (described in Appendix B). We propose to illustrate the effectiveness of our proposed framework in reducing the number of price queries needed to ensure that the worst-case ratio is at least $99\%$ and compare different stopping criteria combined with the Ternary search algorithm.
In particular, we sample random instances from a parametric family of demand curves (either linear demand cures or exponential ones) and, for each instance, report the number of price queries needed to achieve 99% Ratio performance when $i)$ using the common stopping criterion in Ternary Search, $ii)$ using our criterion with $\pazocal{C}=\pazocal{F}$, i.e. when we know that the demand is regular and $iii)$ using our criterion with $\pazocal{C}=\pazocal{G}$, i.e., without using the fact the demand is regular. We set $[\underline{v},\overline{v}]=[1,100]$ (see implementation details in Appendix B). We report in Figure 5 our results.
a) Linear underlying demand.
b) Exponential underlying demand.
Figure 5: Number of price queries needed to achieve 99% Ratio performance.
Figure 5 shows that using our characterization of the worst-case ratio allows to stop considerably earlier than using standard stopping rules with Ternary search. When using our criterion without including the information that the demand is regular, we are able to divide by two the number of price queries needed on average compared to the usual bounds. Furthermore, by incorporating the regularity assumption in our criterion we can reduce by a factor $4$ the query complexity. We also note that our stopping criterion is data-driven and changes as a function of the observed conversion rates: for some instances, we stop after only 5 price queries (as opposed to $44$ with usual rules).
4 Conclusion
We present in this work a framework to characterize the value of data obtained through price experiments. We use this framework both to quantify the value of diverse type of historical data but also to prescribe efficient ways of running price experiments. Throughout this work, we showed that local experiments (such as the sign or the value of the gradient at an offered price) may already contain a lot of information. We also illustrated how our framework can be used to select the next price at which one should experiment and the value associated with such decisions. Finally, our analysis allows to significantly reduce the length of the experiment even when committing to standard dynamic algorithms (such as Ternary search) by providing a data-driven stopping criterion.
We believe that this work opens up many exciting research avenues. One natural question would be to extend our result to be applicable to quantify the value of noisy observations of the demand. Furthermore, the use of our method to prescribe the last price at which to experiment before committing to a revenue maximizing price (as discussed in Section 3.2) provides an interesting first step towards more efficient algorithms trading-off exploration and exploitation in dynamic pricing.
5 Auxiliary Results
Proof of Lemma 1.
The single crossing result for regular distributions mentioned in Lemma 1 follows directly from (Allouah et al., 2022, Lemma 2) with the following additional identity:
$q_{s}\cdot\Gamma\left(\Gamma^{-1}\left(\frac{q_{s^{\prime}}}{q_{s}}\right)\frac{v-s}{s^{\prime}-s}\right)$ $=\frac{1}{\frac{1}{q_{s}}\left(1+\left(\frac{q_{s}}{q_{s^{\prime}}}-1\right)\frac{v-s}{s^{\prime}-s}\right)}$
$=\frac{1}{\frac{1}{q_{s}}+\left(\frac{1}{q_{s^{\prime}}}-\frac{1}{q_{s}}\right)\frac{v-s}{s^{\prime}-s}}=\Gamma\left(\Gamma^{-1}\left(q_{s}\right)+\frac{\Gamma^{-1}\left(q_{s^{\prime}}\right)-\Gamma^{-1}\left(q_{s}\right)}{s^{\prime}-s}(v-s)\right)$
For General distributions, the result follows directly from the non-increasing property of CCDF functions. ∎
Proposition 3 (Feasible information sets).
Let $\mathscr{I}_{\mathbf{p},\mathbf{q}}=\{(p_{i},q_{i}):i=0\ldots,N+1\}$ be a set of sorted increasing prices and associated ccdf values with $(p_{0},q_{0})=(\underline{v},1)$ and $(p_{N+1},q_{N+1})=(\overline{v},0)$. Then:
  • $\pazocal{G}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-empty if and only if the sequence $(q_{i})_{0\leq i\leq N+1}$ is non-increasing.
  • $\pazocal{F}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-empty if and only if the sequence $(q_{i})_{0\leq i\leq N+1}$ is non-increasing and the sequence of slopes $\left(\frac{\Gamma^{-1}\left(q_{i+1}\right)-\Gamma^{-1}\left(q_{i}\right)}{p_{i+1}-p_{i}}\right)_{0\leq i\leq N-1}$ is non-decreasing.
Proof of Proposition 3.
Let $\mathscr{I}_{\mathbf{p},\mathbf{q}}=\{(p_{i},q_{i}):i=0\ldots,N+1\}$ a set of sorted increasing prices and associated ccdf values with $(p_{0},q_{0})=(\underline{v},1)$ and $(p_{N+1},q_{N+1})=(\overline{v},0)$.
Case $\pazocal{C}=\pazocal{F}:$
$\Longrightarrow$) If $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ is $\pazocal{F}$-feasible then there exists a distribution characterized by its cdf $F$ in $\pazocal{F}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Then by definition of a cdf, the sequence $(q_{i}=\overline{F}(p_{i}))_{0\leq i\leq N+1}$ is non-increasing. Fix $11\leq i\leq N-$ 1, Using Lemma 1 on $\overline{F}$ with $(s,q_{s})=(p_{i-1},q_{i-1})$ and $(s^{\prime},q_{s^{\prime}})=(p_{i+1},q_{i+1})$, we have
$\overline{F}(p_{i})=q_{i}\geq\Gamma\left(\Gamma^{-1}\left(q_{i-1}\right)+\frac{\Gamma^{-1}\left(q_{i+1}\right)-\Gamma^{-1}\left(q_{i-1}\right)}{p_{i+1}-p_{i-1}}(p_{i}-p_{i-1})\right)$
which implies that $\frac{{\Gamma^{-1}\left(q_{i}\right)}-{\Gamma^{-1}\left(q_{i-1}\right)}}{{p_{i}}-{p_{i-1}}}\leq\frac{{\Gamma^{-1}\left(q_{i+1}\right)}-{\Gamma^{-1}\left(q_{i-1}\right)}}{{p_{i+1}}-{p_{i-1}}}.$ Furthermore, note that
$\frac{{\Gamma^{-1}\left(q_{i+1}\right)}-{\Gamma^{-1}\left(q_{i-1}\right)}}{{p_{i+1}}-{p_{i-1}}}=\frac{\Gamma^{-1}(\overline{F}(p_{i+1}))-\Gamma^{-1}(\overline{F}(p_{i-1}))}{{p_{i+1}}-{p_{i-1}}}\leq\frac{\Gamma^{-1}(\overline{F}(p_{i+1}))-\Gamma^{-1}(F(p_{i}))}{{p_{i+1}}-p_{i}}=\frac{{\Gamma^{-1}\left(q_{i+1}\right)}-{\Gamma^{-1}\left(q_{i}\right)}}{{p_{i+1}}-{p_{i}}},$
where the inequality holds because $\Gamma^{-1}(\overline{F}(\cdot))$ is convex (see proof of Lemma 2 in Allouah et al. (2022)) and the slope function of a convex function is non-decreasing.
Therefore, $\frac{{\Gamma^{-1}\left(q_{i}\right)}-{\Gamma^{-1}\left(q_{i-1}\right)}}{{p_{i}}-{p_{i-1}}}\leq\frac{{\Gamma^{-1}\left(q_{i+1}\right)}-{\Gamma^{-1}\left(q_{i}\right)}}{{p_{i+1}}-{p_{i}}}$, which proves the implication.
[…]
Case $\pazocal{C}=\pazocal{G}:$
$\Longrightarrow$) If $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ is $\pazocal{G}$-feasible then there exists a distribution characterized by its cdf $F$ in $\pazocal{G}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Then by definition of a cdf, the sequence $(q_{i}=\overline{F}(p_{i}))_{0\leq i\leq N+1}$ is non-increasing.
[…]
Lemma 6.
Fix three scalars $\alpha,\beta>0$ and $s\geq 0$ and a cumulative distribution function $F$ such that $\overline{F}=\Gamma\left(\alpha+\beta(v-s)\right)$ for $v\geq s$. Then $F$ admits a constant virtual value function equal to $s-\frac{1+\alpha}{\beta}.$
Proof of Lemma 6.
Let us first explicitly compute the virtual value function. The derivative of $\Gamma\left(\alpha+\beta(v-s)\right)$, for any $v\geq s$, is equal to $-\beta\Gamma\left(\alpha+\beta(v-s)\right)^{2}.$ Therefore, the virtual value function evaluated at $v\geq s$ is equal to,
$v-\frac{\Gamma\left(\alpha+\beta(v-s)\right)}{\beta\Gamma\left(\alpha+\beta(v-s)\right)^{2}}=v-\frac{1}{\beta\Gamma\left(\alpha+\beta(v-s)\right)}=v-\frac{1+\alpha+\beta(v-s)}{\beta}=s-\frac{1+\alpha}{\beta}.\qed$
Lemma 7.
For any $r^{*}$ such that ${\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-empty, $\sup{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})=\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})$.
Proof of Lemma 7.
Step 1: We first show that, $\sup{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\leq\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Let $q\in{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ and let $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ such that ${r^{*}}$ is an optimal price for $F$ with associated conversion rate $q$. Consider $i\in\{0,\ldots,N\}$ such that ${r^{*}}\in[p_{i},p_{i+1}]$ The crossing lemma implies that, $q=\overline{F}({r^{*}})\leq\min\left(\overline{G}_{\pazocal{F},{i-1}}({r^{*}}),\overline{G}_{\pazocal{F},{i+1}}({r^{*}})\right)\leq\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}}),$ where the last inequality holds by definition of $\overline{U}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ when $\pazocal{C}$ is either $\pazocal{G}$ or $\pazocal{F}$.
Step 2: Next, we show that $\sup{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\geq\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ by proving that $\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})\in{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. To do so, we need to show that there exists a distribution $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ such that ${r^{*}}$ is an optimal price and $\overline{F}({r^{*}})=\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})$. We consider as a candidate the distribution $F_{\pazocal{C}}(\cdot|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}})$ defined in Equation 2.10. Let $\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+}=\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}}))$ and note that, $U_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$ and therefore this set is non-empty. Hence, by Lemma 3, $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$, which implies that, $F_{\pazocal{C}}(\cdot|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}})\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\subset\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$.
To show that ${r^{*}}$ is the optimal price, we see from Lemma 4 that it suffices to show that ${r^{*}}\cdot\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})\geq\max_{i}p_{i}\cdot q_{i}$. Let $q\in{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ then by definition there exists $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ such that $\overline{F}({r^{*}})=q$ and ${r^{*}}\cdot\overline{F}({r^{*}})\geq p\cdot\overline{F}(p)$ for every $p\in[\underline{v},\overline{v}]$. This implies that for every $i\in\{0,N+1\}$,
${r^{*}}\cdot\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})\stackrel{{\scriptstyle(a)}}{{\geq}}{r^{*}}\cdot q={r^{*}}\cdot\overline{F}({r^{*}})\geq p_{i}\cdot\overline{F}(p_{i})=p_{i}\cdot q_{i},$
where $(a)$ follows from step 1. This concludes the proof. ∎
Appendix for
Fast revenue maximization with few experiments
Achraf Bahamou, Columbia University, IEOR department, achraf.bahamou@columbia.edu
Omar Besbes, Columbia University, Graduate School of Business, obesbes@columbia.edu
Omar Mouchtaki, NYU Stern School of Business, om2166@nyu.edu
Appendix A Details on numerical evaluations in Section 3
In Section 3, we provide an understanding of the best achievable Ratio across all pricing mechanisms in $\pazocal{P}$ for different choices of $\mathscr{I}_{\mathbf{p},\mathbf{q}}$.
We next, provide more details about how we numerically derive a lower bound on (Maximin) by leveraging the reduction in Theorem 1. First, we note that the epigraph formulation of (Maximin) satisfies,
$\!\sup_{\Psi(\cdot)\in\pazocal{P},\,\lambda\in[0,1]}$ $\lambda$
subject to $\sup_{F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}R_{\lambda}\left(\Psi,F\right)\leq 0.$
The constraint in the problem involves the infinite-dimensional optimization problem defined in (1.2). Hence, by applying Theorem 1 we obtain the equivalent problem,
$\!\sup_{\Psi(\cdot)\in\pazocal{P},\,\lambda\in[0,1]}$ $\lambda$
subject to $R_{\lambda}\left(\Psi,F_{\pazocal{C}}(\cdot|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)\leq 0\quad\text{ for every }{r^{*}}\in[\underline{v},\overline{v}]\text{ s.t. }{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\neq\emptyset.$
Finally, we derive a lower bound on Problem (Maximin) by restricting the space of randomized mechanisms to discrete mechanisms (i.e., mechanisms such that the distribution of prices a discrete over $[\underline{v},\overline{v}]$) and projecting the one dimensional set of constraints into the associated support of the mechanisms. More specifically, consider an increasing sequence of $M+2$ positive reals $\mathbb{A}=\{a_{i}\}_{i=0}^{M+1}$ a partition of the feasible interval $\mathscr{S}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ for $M\mbox{ in }\mathbb{N}^{*}$ and define the set of discrete mechanisms on $\mathbb{A}$ as
$\pazocal{P}_{\mathbb{A}}=\left\{\Psi\in\pazocal{P}\>:\>\Psi(x)=\sum_{j=0}^{M+1}p_{j}\mathbf{1}\{x\geq a_{i}\},\mbox{ for some }0\leq p_{i}\leq 1,\>\sum_{j=0}^{M+1}p_{i}=1\right\}.$
We will define linear programs parameterized by the finite sequence of increasing positive reals $\mathbb{A}=\{a_{i}\}_{i=0}^{M+1}$ and the information set at hand $\mathscr{I}_{\mathbf{p},\mathbf{q}}$. In particular, we define the following linear programs.
$\underline{\pazocal{L}}(\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}),\mathbb{A})=$ $\!\sup_{\mathbf{p},\lambda}$ $\lambda$
subject to $\lambda\cdot\mbox{opt}\left(\overline{F}_{\pazocal{C}}(\cdot|a_{i+1}^{-},\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)-\sum_{j=0}^{M}Rev\left(a_{j}|\overline{F}_{\pazocal{C}}(\cdot|a_{i},\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)p_{j}\leq 0\quad i=0,...M+1,$
$\sum_{j=0}^{M+1}p_{j}\leq 1,\quad p_{i}\geq 0\quad i=0,...M+1,\quad 0\leq\lambda\leq 1,$
with $\mbox{opt}(F_{\pazocal{C}}(\cdot|a_{i+1}^{-},\mathscr{I}_{\mathbf{p},\mathbf{q}}))=\lim_{x\to a_{i+1}^{-}}\mbox{opt}(F_{\pazocal{C}}(\cdot|x,\mathscr{I}_{\mathbf{p},\mathbf{q}}))$ for any $i=0,...M$.
We note that a lines of technical arguments similar to the one used in (Allouah et al., 2023) enable to show that $\underline{\pazocal{L}}(\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}),\mathbb{A})$ is a lower bound on the value of (Maximin). Furthermore, the solution $(p_{i})_{i\leq N+2}$ provides a feasible distribution of prices. Finally, we note that by judiciously constructing a discrete grid one can ensure that the values associated with the sequence of linear programs converge to the one of (Maximin) as the grid becomes finer.
In all the applications discussed in Section 3, we use the grid of $M=2500$.
Appendix B Details about Section 3.3
We first present formally the Ternary Search procedure in Algorithm 2.
Initialize $a=\underline{v}$, $b=\overline{v}$, $\pazocal{H}_{0}=\{(\underline{v},1),(\overline{v},0)\}$;
for t=1,2,3… do
       Price at $p_{1}=a+\frac{b-a}{3}$, observe $q_{1}$ ;
       Price at $p_{2}=a+\frac{2(b-a)}{3}$, observe $q_{2}$ ;
       Record observed conversion rates $\pazocal{H}_{t+1}=\pazocal{H}_{t}\cup\{(p_{1},q_{1}),(p_{2},q_{2})\}$ ;
      if $p_{1}\cdot q_{1}<p_{2}\cdot q_{2}$ then
             Set $a=p_{1}$ ;
       end if
      else
             Set $b=p_{2}$ ;
       end if
end for
Return $a$ if $a\cdot q_{a}\geq b\cdot q_{b}$ else return $b$
Algorithm 2 Ternary Search
Implementation details for Figure 5. The linear model satisfies for every $v\in[0,\infty)$,
$\overline{F}(v)=\min\left(\max(a-b\cdot v),0),1\right),$
where $b$ is uniformly sampled from $[1,5]$ and $a$ is uniformly sampled from $[1,b]$.
For the exponential demand we have for every $v\in[0,\infty)$,
$\overline{F}(v)=\min(\exp(a-b\cdot v),1),$
where $b$ is uniformly sampled from $[1,5]$ and $a$ is uniformly sampled from $[-0.2,b]$.
Finally, the stopping criterion used to derive the classical bound in Figure 5 satisfies for every desired level of maximin ratio $\epsilon>0$,
$N(\epsilon)=\frac{\log(\frac{\overline{v}-\underline{v}}{\overline{v}\cdot\epsilon})}{\log(3/2)}.$
Indeed, denote by $T_{N}=[a_{N},b_{N}]$, the interval obtained at iteration $N$ in Algorithm 2, by uni-modality of $\mathrm{Rev}\left(\cdot|\overline{F}\right)$ we know that $r_{F}$ is in $[a_{N},b_{N}]$, therefore we have:
$1-\frac{\mathrm{Rev}\left(a_{N}|\overline{F}\right)}{\mathrm{Rev}\left(r_{F}|\overline{F}\right)}$ $=\frac{r_{F}\cdot\overline{F}(r_{F})-a_{N}\cdot\overline{F}(a_{N})}{\mathrm{Rev}\left(r_{F}|\overline{F}\right)}$
$\leq\frac{(r_{F}-a_{N})\cdot\overline{F}(r_{F})}{\mathrm{Rev}\left(r_{F}|\overline{F}\right)}\leq\frac{b_{N}-a_{N}}{\mathrm{Rev}\left(r_{F}|\overline{F}\right)}=\frac{\overline{v}-\underline{v}}{\mathrm{Rev}\left(r_{F}|\overline{F}\right)}\left(\frac{2}{3}\right)^{N}\leq\frac{\overline{v}-\underline{v}}{\underline{v}}\left(\frac{2}{3}\right)^{N}$
Therefore for $N\geq N(\epsilon)=\frac{\log(\frac{\overline{v}-\underline{v}}{\overline{v}\cdot\epsilon})}{\log(3/2)}$, we obtain the desired level of the maximin ratio.