### v2

# Fast Revenue Maximization

Abstract
Problem definition: We study a data-driven pricing problem in which a seller offers a price for a single item based on demand observed at a small number of historical prices. Our goal is to derive precise evaluation procedures of the value of the historical information gathered by the seller, along with prescriptions for more efficient price experimentation. Methodology/results: Our main methodological result is an exact characterization of the maximin ratio (defined as the worst-case revenue garnered by a seller who only relies on past data divided by the optimal revenue achievable with full knowledge of the distribution of values). This result allows to measure the value of any historical data consisting of prices and corresponding conversion rates. We leverage this central reduction to provide new insights about price experimentation. Managerial implications: Motivated by practical constraints that impede the seller from changing prices abruptly, we first illustrate our framework by evaluating the value of local information and show that the mere sign of the gradient of the revenue curve at a single point can provide significant information to the seller. We then showcase how our framework can be used to run efficient price experiments. On the one hand, we develop a method to select the next price experiment that the seller should use to maximize the future robust performance. On the other hand, we demonstrate that our result allows to considerably reduce the number of price experiments needed to reach preset revenue guarantees through dynamic pricing algorithms.
Keywords. pricing, experiment, data-driven algorithms, value of data, distributionally robust optimization, conversion rate, limited information.
1 Introduction and Problem Formulation
In this work, we investigate the monopoly pricing problem in which a decision-maker needs to select a price to sell a single good to a buyer. This operational problem crystallizes the trade-off between the volume of sales and profit margin as setting a price too low hurts the margin but increases the demand for the product whereas setting it too high has the reverse effect.
An important practical challenge in pricing is that the distribution of value of buyers is unknown to the decision-maker. Therefore, decision-makers need to rely on historical information/transaction data, which commonly includes past prices offered and buyers’ purchase decisions at the offered price. Our goal in the present paper is to provide a principled framework for two complementary tasks: i.) quantify the value of classical data structures in the form of prices and associated conversion rates, and ii.) perform efficient price experimentation.
1.1 Problem formulation
We consider the problem of a seller trying to sell one indivisible good to one buyer. We assume that the buyer’s value $v$ is drawn from some unknown distribution $F$ represented by its cumulative distribution function (cdf) and with support included in $[\underline{v},\overline{v}]$, where $0\leq\underline{v}<\overline{v}<\infty$. The seller knows that the distribution of values belongs to some non-parametric class of distributions $\pazocal{C}$. The seller also has information about conversion rates observed at historical prices offered. Formally, let $\mathbf{p}=(p_{i})_{i\in\{0,\ldots N+1\}}$ be a non-decreasing sequence of prices such that $p_{0}=\underline{v}$ and $p_{N+1}=\overline{v}$ and let $\mathbf{q}=(q_{i})_{i\in\{0,\ldots N+1\}}$ be a sequence of historical conversion rates observed such that $q_{0}=1$ and $q_{N+1}=0$. The decision-maker information set is defined as
$\mathscr{I}_{\mathbf{p},\mathbf{q}}=\{(p_{i},q_{i});\;i\in\{0,\ldots,N+1\}\}.$
The problem we are interested in is the following: how can the seller leverage the information contained in $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ to “robustly” maximize her revenue, and how should one guide price experimentation, given that such information has already been collected.
Pricing mechanisms and performance. In this problem, the decision-maker chooses a (potentially) randomized pricing mechanism characterized by the cdf of prices that the seller posts. Formally, we let $\pazocal{P}=\{\Psi\in\pazocal{D}_{[\underline{v},\overline{v}]}\}$ be the set of randomized mechanisms, where $\pazocal{D}_{[\underline{v},\overline{v}]}$ is the set of probability measures on $[\underline{v},\overline{v}]$. Then, for every $\Psi\in\pazocal{P}$ and any distribution of consumer value $F$, the expected revenue generated from the randomized mechanism is defined as,
[…]
where we let $\mathrm{Rev}\left(p|F\right)=p\cdot\mathbb{P}_{v\sim F}\left(v\geq p\right).$ To ease notation, we define the complementary cumulative distribution (ccdf) $\overline{F}(p)=1-F(p)$ for every $p\in\mathbb{R}$ and let $\mathbb{P}_{v\sim F}\left(v\geq p\right)=\overline{F}(p-)$. This notation is consistent with the fact that both $\mathbb{P}_{v\sim F}\left(v\geq p\right)$ and $\overline{F}(p-)$ corresponds to the right limit of $\overline{F}$ at $p$.
In an ideal scenario where the seller has full knowledge of the distribution of values $F$, the optimal revenue is obtained by a posted price mechanism (Riley and Zeckhauser, 1983) and we denote it by $\mbox{opt}(F):=\sup_{p\in[\underline{v},\overline{v}]}\mathrm{Rev}\left(p|F\right).$ We also define the set of optimal prices as $\mathscr{O}(F)=\arg\max_{p\in[\underline{v},\overline{v}]}\mathrm{Rev}\left(p|F\right)$. In this works, we aim at understanding the revenue gap incurred by the seller due to not knowing the actual distribution. Given a mechanism $\Psi\in\pazocal{P}$ and a value distribution $F$, we define this gap through the following ratio,
[…]
Performance for a given information set. Our goal is to understand the spectrum of achievable performances for various information sets $\mathscr{I}_{\mathbf{p},\mathbf{q}}$. In particular, we assume that the seller is given the information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}$. We then define the performance of a pricing mechanism $\Psi\in\pazocal{P}$ as the worst-case ratio, where the worst-case is taken over a class of admissible value distributions denoted as $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ and defined as,
$\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})=\left\{F\in\pazocal{C}\text{ s.t. }\overline{F}(p_{i}-)=q_{i}\text{ for all }(p_{i},q_{i})\in\mathscr{I}_{\mathbf{p},\mathbf{q}}\right\}.$
We note that the set $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ contains all value distributions in $\pazocal{C}$ which are consistent with the information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ observed by the decision-maker.
Consequently, the worst-case performance of any pricing mechanism $\Psi\in\pazocal{P}$ given $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ is the value of the following optimization problem.
$\inf_{F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}\mathrm{Ratio}\left(\Psi,F\right).$ (1.2)
In what follows, we will refer to (1.2) as Nature’s problem. We note that a higher value of this worst-case ratio indicates that the information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ is highly informative in terms of revenue maximization and that the policy $\Psi\in\pazocal{P}$ used by the decision-maker effectively takes advantage of the information in $\mathscr{I}_{\mathbf{p},\mathbf{q}}$.
While (1.2) quantifies the performance of a fixed policy, we study in Section 4 the intrinsic value of certain information sets by computing the best achievable worst-case ratio across all pricing mechanisms. This quantity is formally defined as,
$\mathfrak{R}\left(\pazocal{P},\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\right):=\sup_{\Psi\in\pazocal{P}}\inf_{F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}\mathrm{Ratio}\left(\Psi,F\right).$ (Maximin)
Distribution classes. In this work, we derive results on the performance of pricing mechanisms for two classes of distributions. The class of general distributions denoted as $\pazocal{G}$ that contains all distributions supported on $[\underline{v},\overline{v}]$. We also consider the subset of $\pazocal{G}$ denoted as $\pazocal{F}$ and which denotes the set of regular distributions, i.e., distributions that admit a density, except potentially at the upper boundary of their support and that have a non-decreasing virtual value $v-\frac{1-F(v)}{f(v)}$. We note that the class of regular distributions contains a wide variety of distributions; see, e.g., Ewerhart (2013). This class is central to the pricing and mechanism design literature and can be alternatively described as the class of distributions that induce a concave revenue function in the quantity space.
1.2 Contribution
Performance evaluation for any pricing mechanism. To evaluate the performance of a given pricing mechanism, one needs to solve the optimization problem defined in (1.2). Doing so is a priori daunting as (1.2) is a constrained infinite-dimensional optimization problem over the space $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ which is a non-parametric subset of value distributions. The constraints arise in the form of shape constraints (e.g., when $\pazocal{C}=\pazocal{F}$, distributions have to be regular) or constraints on quantiles of the distribution. Our first main contribution consists in reducing Problem (1.2) to one that is amenable to optimization. Formally, we show the following.
Theorem 1 (Reduction for Policy Evaluation).
Assume $\pazocal{C}$ is equal to $\pazocal{G}$ or $\pazocal{F}$. Fix $N\geq 1$ and let $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ be an information set including $N$ historical prices and conversion rates. Assume $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-empty then, for any mechanism $\Psi$ in $\pazocal{P}$,
$\inf_{F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}\mathrm{Ratio}\left(\Psi,F\right)=\inf_{{r^{*}}\in\mathscr{S}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}\mathrm{Ratio}\left(\Psi,F_{\pazocal{C}}(\cdot|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}})\right),$
where $F_{\pazocal{C}}(\cdot|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is a distribution defined in (2.8) and $\mathscr{S}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is a subset of $[\underline{v},\overline{v}]$ described in Proposition 2.
For any pricing mechanism, Theorem 1 reduces the initial infinite-dimensional optimization problem in (1.2) whose value corresponds to the robust performance of the pricing mechanism, to a one-dimensional problem. This is a critical reduction that can then be used to evaluate and compare pricing-mechanisms under different assumptions about the information the decision-maker has at hand. It also enables to approximate Problem (Maximin) efficiently and derive near-optimal pricing mechanisms for general historical data structures.
We present the proof of Theorem 1 in Section 2. This result significantly generalizes (Allouah et al., 2023, Theorem 1) across several dimensions. First, our reduction also applies when $\pazocal{C}$ is the set of general distributions and therefore allows to investigate settings where the decision-maker does not impose shape constraints. More importantly, our reduction holds for information sets with an arbitrary number of historical prices $N$, while the analysis in Allouah et al. (2023) only holds for $N=1$. This considerably extends the possible information sets $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ that we can investigate and, as we will see, opens up the possibility to explore prescriptions for price experimentation. In turn, it will also allow us to derive novel insights on price experimentation. We also note that while this work emphasizes the ratio objective, our proof is general and readily applies to the absolute regret objective.
New insights on price experimentation. Theorem 1 provides a principled approach to quantify the value of historical demand observations.
We illustrate in Section 4.1 how our method can be used to shed light on the value of certain types of price experiments. We consider the practical setting in which the seller cannot substantially vary the price offered and observe demand realizations at two neighboring prices. We show that our method provides a rigorous way to assess the value of these local experiments and enables to precisely quantify the value of knowing the gradient of the revenue curve at a single price as a function of its magnitude. Quite notably, we show that when the demand is regular, the mere sign of the gradient may already provide considerable information to the seller. More importantly, our procedure allows to perform a rigorous cost-benefit analysis by identifying scenarios in which the sign of the gradient (which can be estimated with cheaper price experiments) already provides sufficient information for better performances, and scenarios for which the numerical value of the gradient (which requires longer and more expensive experiments to be estimated) is important to pin down.
We then use our framework as a way to develop or enrich algorithms for price experimentation. In Section 4.2, we consider a setting in which the seller observed the demand at a single historical point and then may select a new price to experiment with before committing to a future price. We show that by judiciously selecting the second price to experiment with, the seller can significantly increase the worst-case ratio, by $10$ to $20\%$ across all possible demand observed at the first price.
In Section 4.3 we demonstrate that our framework can be used to improve dynamic pricing algorithms. We illustrate that by better quantifying the value of data, one can reduce the extent of price experimentation needed to reach a certain revenue guarantee. We consider the case in which the value distribution is regular (and hence the revenue curve is unimodal), and compare our stopping criterion with the commonly used one when experimenting with the Ternary search algorithm. For example, our refined stopping criterion allows to divide by $2$ to $5$ the number of prices offered to ensure a ratio larger than $99\%$.
1.3 Literature Review
Pricing when facing an unknown value distribution has been considered in a distributionally robust framework with support information (Bergemann and Schlag, 2008; Eren and Maglaras, 2010), with moment information (Azar and Micali, 2013; Carrasco et al., 2018; Chen et al., 2022; Chen, Hu and Wang, 2023; Wang et al., 2024) and with knowledge of quantiles (Azar et al., 2013; Allouah et al., 2023). Variants of this problem were also analyzed in the presence of strategic customers (Caldentey et al., 2016).
The pricing problem has also been studied in the offline data-driven setting where the decision-maker has past data about the value distribution and makes a single pricing decision. In the case where past data takes the form of i.i.d. values sampled from the value distribution, Huang et al. (2015) quantifies the number of samples required to be within a factor $(1-\epsilon)$ from the optimal revenue achievable with full knowledge of the distribution. Fu et al. (2015); Babaioff et al. (2018); Daskalakis and Zampetakis (2020); Allouah et al. (2022) provide performance guarantees when the seller has access to a small number of samples. Chen, Cire, Hu and Lagzi (2023) investigate the performance of data-driven pricing algorithms with transactional data and Biggs (2022) proposes pricing policies based on surrogate loss functions to incorporate contextual information.
In the online setting, Aviv and Pazgal (2002); Araman and Caldentey (2009); Farias and Van Roy (2010) examine dynamic pricing within a Bayesian framework, in which the decision-maker possesses prior knowledge about the parameters of the value distribution. The data structure we consider in this work is related to the one studied in the dynamic pricing and learning literature where the decision-maker observes noisy demand observation at an offered price. Kleinberg and Leighton (2003); Besbes and Zeevi (2009); Broder and Rusmevichientong (2012); Keskin and Zeevi (2014); Singla and Wang (2023) study the performance of dynamic pricing policies trading-off between exploration and exploitation in parametric and non-parametric settings. (Broder, 2011, Chapter 3);Cheung et al. (2017); Chen et al. (2020); Perakis and Singhvi (2023) develop and analyze data-driven pricing policies with limited price changes. Bu et al. (2022) investigates a dynamic pricing problem with linear demand, in which the decision-maker uses both offline and online data. They analyze the value of historical (offline) data by quantifying the achievable regret rate as a function of three statistics of the offline transactional data (size, location, and dispersion). Our work differs from this previous literature along various dimensions: we measure the performance of pricing algorithms with respect to a single-shot pricing decision as opposed to the commonly studied cumulative regret, hence there is no tradeoff between exploration and exploitation in our work. Furthermore, we focus on the setting where the decision-maker observes the exact conversion rate at the offered prices and we derive an exact characterization of the worst-case performance of algorithms that allows to exactly quantify the value of past data. In contrast to Bu et al. (2022), our quantification of the value of data is a function of the whole dataset and therefore enables to provide data-dependent insights. We believe that our results and insights may motivate new algorithmic principles for dynamic pricing policies.
Our work relates to Leme et al. (2023b), which compares the informational value of transactional data and actual observations of each buyer’s value. They show that, asymptotically, observing buyer’s value provides significantly more information for revenue maximization. Our work differs by assuming that the decision-maker has a perfect estimate of the conversion rate at a few number of historical prices and characterizes the exact value of such historical data. It is more closely related to Allouah et al. (2023), which provides a tractable procedure to arbitrarily approximate the maximin ratio when the number of conversion rates $N=1$. Our work extends their results to settings in which the decision-maker has access to demand information across several prices. This allows us to derive new insights and prescriptions regarding efficient price experimentation. In parallel developments, that appeared after we posted an initial draft of this paper, Daei Naby et al. (2024) solve the problem when $N=2$ and when the distribution is assumed to be regular. They provide a closed-form characterization of the optimal deterministic price experiment in the same setting as the one we numerically investigate in Section 4.2. Their focus on deterministic prices, closed forms, and associated insights, complements the understanding of this fundamental problem. (Eren and Maglaras, 2010, Section 4) also analyzes the pricing problem with limited experimentation and characterizes the worst-case ratio for general distributions with information about $N$ conversion rates under the assumption that the value distribution is discrete. Our work generalizes this analysis to arbitrary distributions (both continuous and discrete) and provides a different and unified proof which applies not only to general but also to regular distributions.
Finally, we note that our work broadly relates Chehrazi and Weber (2010) which considers a general robust decision problem with limited measurements and applies their methods in the context of optimal debt-settlement, and to Meister and Nietert (2021); Leme et al. (2023a); Okoroafor et al. (2023) which study the sample complexity required to estimate the cumulative distribution functions with binary feedback. We also contribute to the broad literature which aims at deriving an exact characterization of the value of data in the small to moderate data regime (Schlag, 2006; Stoye, 2009; Besbes and Mouchtaki, 2023; Besbes et al., 2023).
2 Proof of Theorem 1
In this section we prove our main methodological result. Assume $\pazocal{C}$ is equal to $\pazocal{G}$ or $\pazocal{F}$. Fix $N\geq 1$ and let $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ be an information set including $N$ historical prices and conversion rates. Fix a pricing mechanism $\Psi$ in $\pazocal{P}$. We first note that we will prove the result by considering the more general notion of $\lambda$-regret defined for every $\lambda\in[0,1]$ as,
[…]
Taking $\lambda=1$ recovers the absolute regret metric. Furthermore, by using a classical epigraph reformulation (described in Section 3) of the worst-case ratio, one can reduce its analysis to the one of the $\lambda$-regret (see Proposition 1 in Anunrojwong et al. (2023)). In what follows, we prove that the reduction stated in Theorem 1 holds for the $\lambda$-regret, for any $\lambda\in[0,1]$.
The proof generalizes the reasoning developed in (Allouah et al., 2023). In Section 2.1 we construct candidate worst-case distributions for problem (1.2). In Section 2.2, we reduce the infinite-dimensional optimization problem to an optimization problem over two parameters, the optimal price of the worst-case distribution and the conversion rate associated to the optimal price. In Section 2.3 we further simplify the optimization problem by fully characterizing the optimal conversion rate for any fixed optimal price and therefore end up with a one-dimensional optimization problem.
Notation. We define for every $F$ in $\pazocal{D}_{[\underline{v},\overline{v}]}$, the inverse $F^{-1}(q):=\inf\{v\in[\underline{v},\overline{v}]\mbox{ s.t. }F(v)\geq q\}$ for all $q$ in $[0,1]$. For a distribution with positive density function $f$ on its support $[a,b]$, where $0\leq a<\infty$ and $a\leq b\leq\infty$, we denote the virtual value function for $v\in[a,b]$ by $\phi_{F}(v):=v-\frac{\overline{F}(v)}{f(v)}.$ For any $v\geq 0$, we define $\Gamma(v):=\frac{1}{1+v}$ and $\Gamma^{-1}(q):=\frac{1}{q}-1$ denotes the inverse of $\Gamma$.
2.1 Candidate worst-case distributions
We first introduce an important auxiliary distribution which is the central building block of our construction of the candidate worst-case distribution. For any pair of values $(s,s^{\prime})$ such that $0\leq s\leq s^{\prime}$ and $01\geq q_{s}\geq q_{s^{\prime}}>$ 0, we define on $[0,\infty)$ the following ccdfs.
$\overline{G}_{\pazocal{G}}(v|(s,q_{s}),(s^{\prime},q_{s^{\prime}}))$ $=\mathbbm{1}\{v\in[\underline{v},s)\}+\mathbbm{1}\{v\in[s,\overline{v})\}\cdot q_{s^{\prime}}$ (2.1)
$\overline{G}_{\pazocal{F},}(v|(s,q_{s}),(s^{\prime},q_{s^{\prime}}))$ $=\mathbbm{1}\{v\in[\underline{v},\overline{v})\}\cdot\Gamma\left(\Gamma^{-1}\left(q_{s}\right)+\frac{\Gamma^{-1}\left(q_{s^{\prime}}\right)-\Gamma^{-1}\left(q_{s}\right)}{s^{\prime}-s}(v-s)\right).$ (2.2)
These distributions are motivated by the fact that they allow us to derive the following bounds on distributions in $\pazocal{C}$ by deriving a so-called single crossing property.
Lemma 1 (Local Bounds).
Fix $F$ in $\pazocal{C}$ and values $(s,s^{\prime})$ such that $\underline{v}\leq s\leq s^{\prime}$ and $0<q_{s^{\prime}}=\overline{F}(s^{\prime}-)\leq q_{s}=\overline{F}(s-)$. We have that,
$\overline{F}(v)$ $\geq\overline{G}_{\pazocal{C}}(v|(s,q_{s}),(s^{\prime},q_{s^{\prime}}))\qquad\mbox{ if }\>v\in[s,s^{\prime})$
$\overline{F}(v)$ $\leq\overline{G}_{\pazocal{C}}(v|(s,q_{s}),(s^{\prime},q_{s^{\prime}}))\qquad\mbox{ if }\>v\geq s^{\prime}\text{ or }v<s.$
Proof of Lemma 1.
When $\pazocal{C}=\pazocal{F}$, the result follows directly from (Allouah et al., 2022, Lemma 2), which generalizes (Barlow and Proschan, 1975, Chapter 4, Theorem 2.18), with the following additional identity:
$q_{s}\cdot\Gamma\left(\Gamma^{-1}\left(\frac{q_{s^{\prime}}}{q_{s}}\right)\frac{v-s}{s^{\prime}-s}\right)$ $=\frac{1}{\frac{1}{q_{s}}\left(1+\left(\frac{q_{s}}{q_{s^{\prime}}}-1\right)\frac{v-s}{s^{\prime}-s}\right)}$
$=\frac{1}{\frac{1}{q_{s}}+\left(\frac{1}{q_{s^{\prime}}}-\frac{1}{q_{s}}\right)\frac{v-s}{s^{\prime}-s}}=\Gamma\left(\Gamma^{-1}\left(q_{s}\right)+\frac{\Gamma^{-1}\left(q_{s^{\prime}}\right)-\Gamma^{-1}\left(q_{s}\right)}{s^{\prime}-s}(v-s)\right)$
When $\pazocal{C}=\pazocal{G}$, the result follows directly from the non-increasing property of complementary cumulative distribution functions. ∎
Lemma 1 shows that $\overline{G}_{\pazocal{C}}$ provides a way to upper and lower bound any function in the class $\pazocal{C}$ sharing the same conversion rates at $s$ and $s^{\prime}$. Given that all distributions in $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ have the same conversion rate at prices $\mathbf{p}$, we next use Lemma 1 to construct candidate worst-case distributions. In particular, given a feasible information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}=\{(p_{i},q_{i});\;i\in\{0,\ldots,N+1\}\}$ we define, $G_{\pazocal{C},i}(v)={G}_{\pazocal{C}}(v|(p_{i},q_{i}),(p_{i+1},q_{i+1}))$ for every $i\in\{0,\ldots,N\}$. For both the regular and general classes of distributions, we define on $[\underline{v},\overline{v})$:
$\overline{L}_{\pazocal{C}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}})=\overline{G}_{\pazocal{C},i}(v),\quad\mbox{if }v\in[p_{i},p_{i+1}),\mbox{ for }i\in\{0,\ldots,N\}.$ (2.3)
Furthermore, we let
$\overline{U}_{\pazocal{G}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ $=\overline{G}_{\pazocal{G},i-1}(v)=q_{i}\quad\mbox{if }v\in[p_{i},p_{i+1}),\mbox{for }i\in\{0,\ldots,N\}.$
$\overline{U}_{\pazocal{F}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ =$=\begin{cases}\min\{1,\overline{G}_{\pazocal{F},1}(v)\}&\quad\mbox{if }v\in[p_{0},p_{1})\\ \min\{\overline{G}_{\pazocal{F},{i-1}}(v),\overline{G}_{\pazocal{F},{i+1}}(v)\}&\quad\mbox{if }v\in[p_{i},p_{i+1}),\mbox{for }i\in\{1,\ldots,N-2\}\\ \overline{G}_{\pazocal{F},{i-1}}(v)&\quad\mbox{if }v\in[p_{i},p_{i+1}),\mbox{ for }i\in\{N-1,N\}.\end{cases}$
Intuitively, Lemma 1 implies that the ccdf $\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is a lower bound of the ccdf of any distribution in $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Similarly, the ccdf $\overline{U}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ provides an upper bound. We illustrate this in Figure 1. The information set is $\{(p_{i},q_{i}),\,i\in\{1,\ldots,4\}\}=\{(0.47,0.25),(0.70,0.055),(0.89,0.017),(0.98,0.006)\}$ and we set $\underline{v}=0,\,\overline{v}=1$.
a) Bounds on the CCDF values.
b) Bounds on the revenue values.
Figure 1: Illustration of the implications of the bounds derived in Lemma 1.
Figure 1-a depicts in red the information set. The shaded region in green (resp. in blue) depicts the region where the ccdf of regular (resp. general) distributions that are consistent with the information set could fall. Figure 1-b depicts the corresponding region for the revenue curve.
2.2 Reduction to a two-dimensional problem
In this section, we prove our first reduction of Nature’s problem to a two-dimensional optimization problem parameterized by the set of possible values that the optimal price and corresponding conversion rate can take. To that end, let ${\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ denote the set of feasible optimal pairs, i.e.,
${\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}}):=\left\{({r^{*}},{q^{*}})\in[\underline{v},\overline{v}]\times[0,1]:\mbox{there exists }\>F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\mbox{ s.t. }{r^{*}}\in\mathscr{O}(F)\mbox{ and }\overline{F}({r^{*}}-)={q^{*}}\right\}.$
Given such a definition, we have for every $\lambda\in[0,1]$,
$\sup_{F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}R_{\lambda}\left(\Psi,F\right)=\sup_{({r^{*}},{q^{*}})\in{\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}\left\{\lambda\cdot{r^{*}}\cdot{q^{*}}-\inf_{\begin{subarray}{c}F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}):\\ {r^{*}}\in\mathscr{O}(F),\overline{F}({r^{*}}-)={q^{*}}\end{subarray}}\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|F\right)]\right\}$ (2.4)
The above reformulation allows to “decouple” the problem as the first term is fully controlled by ${r^{*}}$ and ${q^{*}}$. The second term, which represents the expected revenue of the pricing mechanism, can then be optimized independently. Our next structural result solves exactly the inner minimization problem.
Proposition 1 (Worst-case Revenue).
Assume $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is not empty. For every $({r^{*}},{q^{*}})\in{\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ and every pricing mechanism $\Psi$, we have that,
$\inf_{\begin{subarray}{c}F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}):\\ {r^{*}}\in\mathscr{O}(F),\overline{F}({r^{*}}-)={q^{*}}\end{subarray}}\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|F\right)]=\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))\right)].$
We note that Proposition 1 and (2.4) imply that Nature’s problem can be written as
$\sup_{F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}R_{\lambda}\left(\Psi,F\right)=\sup_{({r^{*}},{q^{*}})\in{\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}\mathbb{E}_{p\sim\Psi}\left[\lambda\cdot{r^{*}}\cdot{q^{*}}-\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))\right)\right].$ (2.5)
In other words, given ${r^{*}}\in\mathscr{O}(F)$ and ${q^{*}}$ s.t. ${q^{*}}=\overline{F}({r^{*}}-)$, one can pin down the worst-case distribution and associated revenue curve. In what follows, we prove Proposition 1.
Our first lemma shows that our candidate worst-case distribution provides a lower bound on the revenue associated with any value distribution $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$.
Lemma 2.
Consider an information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}$, then for any pricing mechanism $\Psi\in\pazocal{P}$ and any value distribution $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$, we have that, $\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|F\right)]\geq\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)]$, where the ccdf of $L_{\pazocal{C}}$ is defined in (2.3).
Proof of Lemma 2.
Fix $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ and consider $v\in[\underline{v},\overline{v}]$. We first show that, $\overline{F}(v)\geq\overline{L}_{\pazocal{C}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}}).$ Consider $i\in\{0,\ldots,N\}$ such that, $v\in[p_{i},p_{i+1})$. We remark that $\overline{F}(p_{i}-)=q_{i}$ and $\overline{F}_{0}(p_{i+1}-)=q_{i+1}$ because $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Therefore Lemma 1 implies that,
$\overline{F}(v)\geq\overline{G}_{\pazocal{C}}(v|(p_{i},q_{i}),(p_{i+1},q_{i+1}))=\overline{L}_{\pazocal{C}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}}),$
where the last equality holds by definition of $\overline{L}_{\pazocal{C}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Hence, $\overline{F}(v)\geq\overline{L}_{\pazocal{C}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Furthermore, for any pricing mechanism $\Psi\in\pazocal{P}$, we have that,
$\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|F\right)]=\int_{0}^{\infty}p\cdot\overline{F}(p-)\,d\Psi(p)\stackrel{{\scriptstyle(a)}}{{\geq}}\int_{0}^{\infty}p\cdot\overline{L}_{\pazocal{C}}(p-|\mathscr{I}_{\mathbf{p},\mathbf{q}})\,d\Psi(p)=\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)],$
where $(a)$ holds because $\overline{F}(v)\geq\overline{L}_{\pazocal{C}}(v|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ for every $v\in[\underline{v},\overline{v}]$ and therefore by taking a right limit $\overline{F}(v-)\geq\overline{L}_{\pazocal{C}}(v-|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ for every $v\in[\underline{v},\overline{v}]$. This concludes the proof. ∎
Next, we show that our candidate worst-case distribution belongs to $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$.
Lemma 3.
If $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-empty, then $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$.
Proof of Lemma 3.
First, one can verify that for every $i\in\{0,\ldots,N+1\}$, $\overline{L}_{\pazocal{C}}(p_{i}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})=q_{i}.$ Therefore, we need to show that, $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\in\pazocal{C}.$
When $\pazocal{C}=\pazocal{G}$, we note that $L_{\pazocal{G}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-decreasing because $(q_{i})_{i\leq N+1}$ is non-decreasing. Indeed, $(q_{i})_{i\leq N+1}$ because the set $\pazocal{G}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-empty. Hence, $L_{\pazocal{G}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\in\pazocal{G}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$.
When $\pazocal{C}=\pazocal{F}$, we show in Proposition 3 (stated and proved in Section 6) that if $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-empty then $\left(\frac{\Gamma^{-1}\left(q_{i+1}\right)-\Gamma^{-1}\left(q_{i}\right)}{p_{i+1}-p_{i}}\right)_{0\leq i\leq N-1}$ is non-decreasing. We next show that this implies that $\overline{L}_{\pazocal{F}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is regular. Lemma 6 implies that the associated virtual value function satisfies
$\phi_{L_{\pazocal{F}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})}(v)={p_{i}}-({p_{i+1}}-{p_{i}})\frac{1+{\Gamma^{-1}\left(q_{i}\right)}}{{\Gamma^{-1}\left(q_{i+1}\right)}-{\Gamma^{-1}\left(q_{i}\right)}},\quad\mbox{if }v\in[p_{i},p_{i+1}),\mbox{ for }i=0\cdots N$
Hence, the virtual value function is piece-wise constant. Now we need to show that $\phi_{L}(v)$ is non-decreasing. Fix $11\leq i\leq N-$ 1. we evaluate the difference between the two consecutive constant values that the virtual value function is and note that it is equal to,
$\left(1+{\Gamma^{-1}\left(q_{i}\right)}\right)\left[\frac{{p_{i}}-{p_{i-1}}}{{\Gamma^{-1}\left(q_{i}\right)}-{\Gamma^{-1}\left(q_{i-1}\right)}}-\frac{{p_{i+1}}-{p_{i}}}{{\Gamma^{-1}\left(q_{i+1}\right)}-{\Gamma^{-1}\left(q_{i}\right)}}\right]\stackrel{{\scriptstyle(a)}}{{\geq}}0,$
where inequality $(a)$ holds because the sequence $\left(\frac{\Gamma^{-1}\left(q_{i+1}\right)-\Gamma^{-1}\left(q_{i}\right)}{p_{i+1}-p_{i}}\right)_{0\leq i\leq N-1}$ is non-decreasing.
Hence, the virtual value function of $\overline{L}_{\pazocal{F}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-decreasing and $L_{\pazocal{F}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\in\pazocal{F}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. ∎
The final structural result needed in the proof of Proposition 1 characterizes the optimal revenue (and associated conversion rate) for the distribution $\overline{L}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})$.
Lemma 4.
If $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-empty, then $\sup_{p\in[\underline{v},\overline{v}]}\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)=\max_{i\in\{0,\ldots,N+1\}}p_{i}\cdot q_{i}.$
Proof of Lemma 4.
We first note that Lemma 3 implies that $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ belongs to $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Therefore, for every $i\in\{1,\ldots,N+1\}$, we have $\mathrm{Rev}\left(p_{i}|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)=p_{i}\cdot\overline{L}_{\pazocal{C}}(p_{i}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})=p_{i}\cdot q_{i}$. Hence, $\sup_{p\in[\underline{v},\overline{v}]}\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)\geq\max_{i\in\{0,\ldots,N+1\}}p_{i}\cdot q_{i}.$
To prove the reverse inequality, it suffices to show that the revenue function $p\mapsto\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)$ is monotonic on every interval $[p_{i},p_{i+1})$ for $i\in\{0,\ldots,N\}$. By the definition in (2.3), we need to analyze the monotonicity of the function $\kappa:p\mapsto p\cdot\overline{G}_{\pazocal{C},i}(p)$.
When $\pazocal{C}=\pazocal{G}$, we have that $\overline{G}_{\pazocal{C},i}(p)$ is constant on $[p_{i},p_{i+1})$. Hence, $\kappa$ is a linear function and is monotonic on $[p_{i},p_{i+1})$.
When $\pazocal{C}=\pazocal{F}$, we define for every $q\in[\overline{G}_{\pazocal{C},i}(p_{i}),\overline{G}_{\pazocal{C},i}(p_{i+1}))$ the function $\tilde{\kappa}(q)=\overline{G}_{\pazocal{C},i}^{-1}(q)\cdot q$. We note that for every $p\in[p_{i},p_{i+1})$ we have $\kappa(p)=\tilde{\kappa}(\overline{G}_{\pazocal{C},i}(p))$ and, by composition, if $\tilde{\kappa}$ is monotonic then $\kappa$ is monotonic. To show that $\tilde{\kappa}$ is monotonic we note that it corresponds to the revenue function in the quantile space. Its derivative satisfies, $\frac{d}{dq}\tilde{\kappa}(q)=\phi_{G_{\pazocal{C},i}}(\overline{G}_{\pazocal{C},i}^{-1}(q))$, where $\phi_{F}$ is the virtual value function of a distribution $F$. To conclude the proof, we show in Lemma 6 that the virtual value function of $G_{\pazocal{C},i}$ is constant. Therefore, $\tilde{\kappa}$ is linear and monotonic on $[q_{i},q_{i+1})$. ∎
We next combine these lemmas to formally prove Proposition 1.
Proof of Proposition 1.
Consider $({r^{*}},{q^{*}})\in{\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ and consider the extended information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+}:=\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup\{({r^{*}},{q^{*}})\}$. We note that $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$ is the subset of $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ where distributions must also satisfy $\overline{F}({r^{*}}-)={q^{*}}$. We also let $\mathscr{A}=\{F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\text{ s.t. }{r^{*}}\in\mathscr{O}(F)\text{ and }\overline{F}({r^{*}}-)={q^{*}}\}.$ We first show that,
$\inf_{F\in\mathscr{A}}\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|F\right)]\geq\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\right)].$ (2.6)
Indeed by definition, we have that $\mathscr{A}\subset\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$. Therefore, (2.6) follows from Lemma 2 applied to the information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+}$. Next, we show that
$\inf_{F\in\mathscr{A}}\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|F\right)]\leq\mathbb{E}_{\Psi}[\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\right)].$ (2.7)
To do so, we will show that $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\in\mathscr{A}$ by proving that $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ and that $({r^{*}},{q^{*}})$ are optimal prices and associated conversion rates for $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$.
We first show that $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$ is non-empty. Note that since $({r^{*}},{q^{*}})\in{\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ there exists $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ such that $\overline{F}({r^{*}}-)={q^{*}}$. Such a distribution belongs to $\mathscr{A}$ and therefore belongs to $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$. Therefore, this set is not empty. It then follows from Lemma 3 that $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\subset\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$.
In addition, Lemma 4 applied to the information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+}$ implies that,
$\sup_{p\in[\underline{v},\overline{v}]}\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\right)=\max\left(\max_{i\in\{0,\ldots,N+1\}}p_{i}\cdot q_{i},{r^{*}}\cdot{q^{*}}\right).$
Furthermore, we argue that ${r^{*}}\cdot{q^{*}}\geq\max_{i\in\{0,\ldots,N+1\}}p_{i}\cdot q_{i}$. This inequality holds because $\mathscr{A}$ is non-empty and for any distribution $F\in\mathscr{A}$, we have for every $i\in\{0,\ldots,N+1\}$ that ${r^{*}}\cdot{q^{*}}={r^{*}}\cdot\overline{F}({r^{*}}-)\stackrel{{\scriptstyle(a)}}{{\geq}}p_{i}\cdot\overline{F}(p_{i}-)=p_{i}\cdot q_{i},$ where $(a)$ holds because $F\in\mathscr{A}$ and hence ${r^{*}}$ is an optimal price for this distribution. This implies that $\sup_{p\in[\underline{v},\overline{v}]}\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\right)={r^{*}}\cdot{q^{*}}$ and given that $\overline{L}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})={q^{*}}$, it concludes that ${r^{*}}$ is an optimal price for $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$ with associated conversion rate ${q^{*}}$.
Therefore, $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\in\mathscr{A}$ and (2.7) holds. Finally, (2.6) and (2.7) imply Proposition 1. ∎
2.3 Reduction to a one-dimensional problem
In a second step, we further reduce the minimization problem stated in Equation (2.5) to a one-dimensional minimization problem by deriving for every value of ${r^{*}}$ the conversion-rate ${q^{*}}$ which maximizes the $\lambda$-regret. For any ${r^{*}}$, we define ${\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}}):=\{q^{*}\in[0,1]\text{ s.t. }(r^{*},q^{*})\in{\mathscr{}B}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\}.$ Furthermore, we fix $\lambda\in[\mathfrak{R}\left(\pazocal{P},\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\right),1]$ and define the mapping,
$\rho:\begin{cases}[0,1]\to\mathbb{R}_{+}\\ {q^{*}}\mapsto\mathbb{E}_{p\sim\Psi}\left[\lambda\cdot{r^{*}}\cdot{q^{*}}-\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))\right)\right].\end{cases}$
We next show that $\rho$ is monotonic.
Lemma 5.
Fix $\lambda\in[\mathfrak{R}\left(\pazocal{P},\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\right),1]$. Then for every $r^{*}$ such that, ${\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-empty, the mapping $\rho$ is non-decreasing.
Proof of Lemma 5.
We note that for every $q^{*}>0$, $\rho(q^{*})=q^{*}\cdot\frac{\rho(q^{*})}{q^{*}}$. To conclude the proof it is sufficient to show that $\frac{\rho(q^{*})}{q^{*}}$ is non-negative and non-decreasing as the product of two non-negative and non-decreasing functions is non-decreasing.
First, we remark from the epigraph formulation in Section 3 that $\frac{\rho(q^{*})}{q^{*}}$ is non-negative. To prove that $\frac{\rho(q^{*})}{q^{*}}$ is non-decreasing remark that, $\frac{\rho({q^{*}})}{{q^{*}}}=\lambda r^{*}-\mathbb{E}_{p\sim\Psi}\left[\frac{1}{q^{*}}\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))\right)\right]$. Hence, it is sufficient to show that for every $p$, the mapping $\tilde{\rho}:{q^{*}}\mapsto\frac{1}{{q^{*}}}\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))\right)$ is non-increasing.
When $\pazocal{C}=\pazocal{G}$, $\tilde{\rho}$ is a constant when for every $p$ such that $\overline{L}_{\pazocal{C}}(p|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))={q^{*}}$ and is the inverse function when $\overline{L}_{\pazocal{C}}(p|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}}))=q_{i}\neq{q^{*}}$. In both cases, it is non-increasing.
When $\pazocal{C}=\pazocal{F}$, let $\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+}=\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},{q^{*}})$ be the extended information set and reindex the elements such that $\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+}=\{(\hat{p}_{i},\hat{q}_{i});\;i=0\ldots N+2\}$. Fix $p$ and let $i$ be such that $p\in[\hat{p}_{i},\hat{p}_{i+1})$ By definition of $L_{\pazocal{C}}$ we have that
$\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\right)=p\cdot\overline{G}_{\pazocal{C}}(p-|(\hat{p}_{i},\hat{q}_{i}),(\hat{p}_{i+1},\hat{q}_{i+1})).$
Hence, $\mathrm{Rev}\left(p|L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\right)$ only varies as a function of ${q^{*}}$ if ${r^{*}}\in\{\hat{p}_{i},\hat{p}_{i+1}\}$. In all other cases, the numerator in $\tilde{\rho}$ is a constant which does not depend on ${q^{*}}$ and therefore $\tilde{\rho}$ is clearly non-increasing. We next show that when either $p_{i}={r^{*}}$ or $p_{i+1}={r^{*}}$, $\tilde{\rho}$ is also non-increasing.
Case 1: $p_{i}={r^{*}}$. In this case, we remark using the alternative form of $\bar{G}_{\pazocal{C}}$ derived in the proof of Lemma 1 that,
$\tilde{\rho}({q^{*}})=\frac{1}{{q^{*}}}\cdot p\cdot{q^{*}}\cdot\Gamma\left(\Gamma^{-1}\left(\frac{q_{i+1}}{{q^{*}}}\right)\frac{p-{r^{*}}}{p_{i+1}-{r^{*}}}\right)=p\cdot\Gamma\left(\Gamma^{-1}\left(\frac{q_{i+1}}{{q^{*}}}\right)\frac{v-{r^{*}}}{p_{i+1}-{r^{*}}}\right).$
By composition, we conclude that $\tilde{\rho}$ is non-increasing.
Case 2: $p_{i+1}={r^{*}}$. Using again one of the alternative forms of $\overline{G}_{\pazocal{C}}$ derived in the proof of Lemma 1 we obtain that,
$\tilde{\rho}({q^{*}})=\frac{1}{{q^{*}}}\cdot\frac{p}{\frac{1}{q_{i}}+\left(\frac{1}{{q^{*}}}-\frac{1}{q_{i}}\right)\frac{p-p_{i}}{{r^{*}}-p_{i}}}=\frac{p}{\frac{{q^{*}}}{q_{i}}\cdot\left(1-\frac{p-p_{i}}{{r^{*}}-p_{i}}\right)+\frac{p-p_{i}}{{r^{*}}-p_{i}}}.$
By composition, we conclude again that $\tilde{\rho}$ is non-increasing. ∎
To complete the reduction to a one-dimensional problem, we remark that,
$\sup_{F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}R_{\lambda}\left(\Psi,F\right)=\sup_{r^{*}\in[\underline{v},\overline{v}]}\sup_{q^{*}\in{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}\rho(q^{*})$ =$\stackrel{{\scriptstyle(a)}}{{=}}\sup_{r^{*}\in[\underline{v},\overline{v}]}\mathbbm{1}\{{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\neq\emptyset\}\cdot\rho(\sup{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}}))$
=$\stackrel{{\scriptstyle(b)}}{{=}}\sup_{r^{*}\in[\underline{v},\overline{v}]}\mathbbm{1}\{{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\neq\emptyset\}\cdot\rho(\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})).$
Where $(a)$ follows from Lemma 5 and $(b)$ follows from Lemma 7, stated and proved in Section 6.
For every ${r^{*}}\in[\underline{v},\overline{v}]$, we consider the function defined for every $v\in[\underline{v},\overline{v}]$ as,
$\overline{F}_{\pazocal{C}}(v|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}}):=\overline{L}_{\pazocal{C}}\left(v|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}}))\right).$ (2.8)
We note that we have established that,
$\sup_{F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}R_{\lambda}\left(\Psi,F\right)=\sup_{{r^{*}}\in[\underline{v},\overline{v}]}\mathbbm{1}\{{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\neq\emptyset\}\cdot R_{\lambda}\left(\Psi,F_{\pazocal{C}}(\cdot|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)=\sup_{{r^{*}}\in\mathscr{S}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}R_{\lambda}\left(\Psi,F_{\pazocal{C}}(\cdot|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}})\right),$
where the last equality holds by defining $\mathscr{S}(\mathscr{I}_{\mathbf{p},\mathbf{q}})=\{{r^{*}}\in[\underline{v},\overline{v}]\text{ s.t. }{\mathscr{}B}_{{r^{*}}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\neq\emptyset\}$. This concludes our reduction to a one-dimensional optimization problem and concludes the proof of Theorem 1.
Evaluation of the condition: ${\mathscr{}B}_{{r^{*}}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\neq\emptyset$. When performing a line search by using the reduction we developed, one needs to assess whether ${r^{*}}\in\mathscr{S}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Our next result shows that this condition can be verified efficiently.
Proposition 2.
Fix ${r^{*}}\in[\underline{v},\overline{v}]$. Then ${r^{*}}\in\mathscr{S}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ if and only if,
$i)$ $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})))$ is non-empty, and $ii)$ ${r^{*}}\cdot\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})\geq\max_{i\in\{0,\ldots,N+1\}}p_{i}\cdot q_{i}$.
Proof of Proposition 2.
First, if ${\mathscr{}B}_{{r^{*}}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$, Lemma 7 (in Section 6) implies that $\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})\in{\mathscr{}B}_{{r^{*}}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. By definition, of ${\mathscr{}B}_{{r^{*}}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ this implies that $i)$ and $ii)$ are satisfied.
Reciprocally, if $\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})))$ is non-empty, then Lemma 3 implies that the distribution $\overline{L}_{\pazocal{C}}\left(v|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}}))\right)\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})))\subset\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Furthermore, Lemma 4 along with the condition that ${r^{*}}\cdot\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})\geq\max_{i\in\{0,\ldots,N+1\}}p_{i}\cdot q_{i}$, implies that ${r^{*}}$ is an optimal price for $\overline{L}_{\pazocal{C}}\left(v|\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}}))\right)$ with associated conversion-rate $\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})$. This proves that $\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})\in{\mathscr{}B}_{{r^{*}}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Hence, this set is non-empty. ∎
3 Quantifying the Value of Information
Theorem 1 implies that one can efficiently quantify the worst-case performance of any pricing mechanism given a particular information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}$. We next, establish how this allows to derive a lower bound on (Maximin). We note that (Maximin) characterizes the intrinsic value of an information set $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ as this ratio represents the best achievable performance across all pricing mechanisms having access to past prices and conversion rates in $\mathscr{I}_{\mathbf{p},\mathbf{q}}$. Evaluating (Maximin) will be critical to derive new insights on price experimentation in Section 4.
The following discussion presents an application of the method developed in (Allouah et al., 2023) to our setting. First, we note that the epigraph formulation of (Maximin) satisfies,
$\!\sup_{\Psi(\cdot)\in\pazocal{P},\,\lambda\in[0,1]}$ $\lambda$
subject to $\sup_{F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})}R_{\lambda}\left(\Psi,F\right)\leq 0.$
The constraint in the problem involves the infinite-dimensional optimization problem defined in (1.2). Hence, by applying Theorem 1 we obtain the equivalent problem,
$\!\sup_{\Psi(\cdot)\in\pazocal{P},\,\lambda\in[0,1]}$ $\lambda$
subject to $R_{\lambda}\left(\Psi,F_{\pazocal{C}}(\cdot|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)\leq 0\quad\text{ for every }{r^{*}}\in\mathscr{S}(\mathscr{I}_{\mathbf{p},\mathbf{q}}).$
Finally, we derive a lower bound on (Maximin) by restricting the space of randomized mechanisms to discrete mechanisms (i.e., mechanisms such that the distribution of prices is discrete over $[\underline{v},\overline{v}]$) and by projecting the one dimensional set of constraints into the associated support of the mechanisms. More specifically, consider an increasing sequence of $M+2$ positive reals $\mathbb{A}=\{a_{i}\}_{i=0}^{M+1}$ forming a partition of the feasible interval $\mathscr{S}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ for some $M\in\mathbb{N}^{*}$ and define the set of discrete mechanisms on $\mathbb{A}$ as
$\pazocal{P}_{\mathbb{A}}=\left\{\Psi\in\pazocal{P}\>:\>\Psi(x)=\sum_{i=0}^{M+1}\psi_{i}\cdot\mathbf{1}\{x\geq a_{i}\},\mbox{ for some }0\leq\psi_{i}\leq 1,\>\sum_{j=0}^{M+1}\psi_{i}=1\right\}.$
We will define linear programs parameterized by the finite sequence of increasing positive reals $\mathbb{A}=\{a_{i}\}_{i=0}^{M+1}$ and the information set at hand $\mathscr{I}_{\mathbf{p},\mathbf{q}}$. In particular, we define the following linear programs.
$\underline{\pazocal{L}}(\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}),\mathbb{A})=\;$ $\!\sup_{\mathbf{p},\lambda}$ $\lambda$ (3.3a)
s.t. $\lambda\cdot\mbox{opt}\left(F_{\pazocal{C}}(\cdot|a_{i+1}^{-},\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)-\sum_{j=0}^{M}\mathrm{Rev}\left(a_{j}|F_{\pazocal{C}}(\cdot|a_{i},\mathscr{I}_{\mathbf{p},\mathbf{q}})\right)\psi_{j}\leq 0\quad i\in\{0,\ldots,M+1\},$ (3.3b)
$\sum_{j=0}^{M+1}\psi_{j}\leq 1,\quad\psi_{i}\geq 0\quad i\in\{0,\ldots,M+1\},\quad 0\leq\lambda\leq 1,$ (3.3c)
where $\mbox{opt}(F_{\pazocal{C}}(\cdot|a_{i+1}^{-},\mathscr{I}_{\mathbf{p},\mathbf{q}}))=\lim_{x\to a_{i+1}^{-}}\mbox{opt}(F_{\pazocal{C}}(\cdot|x,\mathscr{I}_{\mathbf{p},\mathbf{q}}))$ for any $i=0,...M$.
We note that a line of technical arguments similar to the one used in (Allouah et al., 2023) enables to show that $\underline{\pazocal{L}}(\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}),\mathbb{A})$ is a lower bound on the value of (Maximin). Furthermore, the solution $(\psi_{i})_{i\leq N+2}$ provides a feasible distribution of prices. Finally, we note that by judiciously constructing a discrete grid one can ensure that the values associated with the sequence of linear programs converge to the one of (Maximin) as the grid becomes finer.
4 Applications to Price Experimentation
We described in Section 3 a tractable procedure to quantify the value of any historical information set that comes in the form of prices and their associated conversion rates. In this section, we illustrate how our analysis can be leveraged to derive insights on price experimentation. In Section 4.1, we study the value of local experimentation and consider a setting in which the decision-maker cannot substantially vary prices and therefore has only an estimate of the gradient at a given price along with the conversion rate at that price. In Section 4.2, we investigate the value of global experimentation and derive an optimal (freely selected) second price to experiment at after receiving the conversion rate from a first price. Finally, in Section 4.3, we illustrate how our exact analysis could be used to considerably reduce the number of price experiments needed to reach a certain level of performance guarantee.
4.1 Pricing with gradient information: the value of local experimentation
In many practical settings, the seller cannot drastically change the offered price and therefore can only gather “local” information. Motivated by this, we consider the setting in which the seller knows the support of the demand function $[\underline{v},\overline{v}]$ and has observed two demand points, i.e. $\mathscr{I}_{\mathbf{p},\mathbf{q}}=\{(p_{1},q_{1}),(p_{\epsilon},q_{\epsilon})\}$. The two prices are such that $p_{\epsilon}=(1+\epsilon)p_{1}$ for $\epsilon>0$.
Value of the numerical gradient. We propose to understand the value of such observations by leveraging the procedure described in Section 3 to numerically compute the maximin ratio defined in (Maximin) when the demand function is regular (i.e., $\pazocal{C}=\pazocal{F}$) and its support is included in $[\underline{v},\overline{v}]=[1,100]$. We consider the setting in which the seller experiments at an increment of $\epsilon=1\%$ of the observed price $p_{1}$ which we fix at $10$ and we consider different initial observed demand $q_{1}$ in $\{0.01,0.1,0.25,0.5,0.75,0.9,0.99\}$. For each initial $q_{1}$, we evaluate a lower bound on the maximin ratio for a grid of 200 possible values of $q_{\epsilon}$ and report the results in Figure 2. For the sake of interpretability, the x-axis of Figure 2 is parametrized by $g_{\epsilon}$ which is defined as,
$g_{\epsilon}=\frac{p_{\epsilon}q_{\epsilon}-p_{1}q_{1}}{p_{\epsilon}-p_{1}}.$
We interpret $g_{\epsilon}$ as the gradient of the revenue curve at $p_{1}$ given that, $\nabla_{p}\mathrm{Rev}(p_{1}|F)\approx\frac{p_{\epsilon}q_{\epsilon}-p_{1}q_{1}}{p_{\epsilon}-p_{1}}$.
a) Negative observed gradient.
b) Positive observed gradient.
Figure 2: Value of a gradient measurement for various $q_{1}$ ($\pazocal{C}=\pazocal{F}$). The maximin ratio is computed by evaluating $\underline{\pazocal{L}}(\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}),\mathbb{A})$ with a discretization involving $M=2500$ points.
Figure 2 naturally demonstrates that when the absolute value of the gradient is low, the worst-case ratio is guaranteed to be high as the decision-maker is very close to the optimal price. Interestingly, relatively large values of the gradient are still very informative and allow considerable improvement compared to a single price observation. When $q_{1}=0.5$, i.e., when the first price was at the median of the value distribution, the worst-case ratio without gradient information is around $60\%$ (as we will see in Figure 3). By knowing that the absolute value of the gradient is lower or equal to $0.1$, the worst-case ratio can increase to more that $80\%$.
Value of the gradient-sign: In settings where the estimation of the absolute value of the gradient is noisy and unreliable and where the seller can only estimate the sign of the gradient, our framework allows to characterize the value of such measurement when the demand is regular as it translates to the elimination of one side of the interval of feasible optimal reserve prices. Indeed, given that the revenue function is unimodal in that case, the sign of the gradient at a price $p_{1}$ allows to assess whether the optimal price ${r^{*}}$ satisfies ${r^{*}}\geq p_{1}$ or ${r^{*}}\leq p_{1}$.
We next illustrate numerically the value of a positive and negative gradient measurement in the setting in which the seller knows that the demand function is regular and its support is included in $[\underline{v},\overline{v}]=[1,100]$. As a baseline, we first consider the maximin ratio obtained by the retailer when the information set is $\{(p_{1},q_{1})\}$, i.e., when the seller knows the conversion rate at a single price. We then evaluate the maximin ratio obtained by the seller who also knows the gradient sign at the observed price $p_{1}$. For each initial conversion rate $q_{1}$, we evaluate the maximin ratio for negative and positive observed gradient-signs as well as the value of the maximin ratio without gradient information (our baseline) and report the values in Figure 3.
Figure 3: Value of the gradient-sign for various $q_{1}$. The maximin ratio with gradient-sign information is computed by modifying the evaluation of $\underline{\pazocal{L}}(\pazocal{C}(\{(p_{1},q_{1})\}),\mathbb{A})$ where we only consider the constraints in (3.3b) for $i$ such that $a_{i}\leq p_{1}$ (resp. $a_{i}\geq p_{1}$) when the gradient-sign is negative (resp. positive). ($p_{1}=10$)
[…]
This example shows that depending on the initial conversion rate $q_{1}$ the marginal value associated with knowing the sign of the gradient or the numerical value of the gradient varies considerably. Our procedure allows to quantify the value of running longer and more accurate (hence more expensive) local price experiments as a function of the data at hand.
4.2 The value of global experimentation
We now relax the constraint that the decision-maker should explore locally and consider the setting in which the seller has observed a single demand point $(p_{1},q_{1})$ and has the opportunity to experiment at any second price $p_{2}$ by collecting the demand level at the latter price, before committing to a final pricing mechanism. We model the problem as a game between Nature and the seller, in which the seller commits to a price $p_{2}$ to experiment at, and Nature counters the seller by choosing the associated feasible conversion rate $q_{2}$ to yield the worst maximin ratio in the final stage. More specifically, for a fixed $p_{2}$, Nature is solving the following problem,
$\inf_{q_{2}\in[0,1]}\sup_{\Psi\in\pazocal{P}}\inf_{F\in\pazocal{C}(\{(p_{1},p_{2}),(q_{1},q_{2})\})}\mathrm{Ratio}\left(\Psi,F\right).$
We illustrate numerically the value of one additional experiment in the setting where the seller knows the demand function is regular and its support is included in $[\underline{v},\overline{v}]=[1,100]$. The seller starts with one initial observed demand at the price $p_{1}=10$. For each initial $q_{1}$ and candidate experiment price $p_{2}$ from a fixed grid of $[\underline{v},\overline{v}]$ of size 200, we numerically evaluate a lower-bound on the above objective using the LP described in Section 3 (with $M=2500$). We present in Figure 4-a, the improvement in the worst-case ratio allowed with the best experiment price $p_{2}$ chosen by the decision-maker and show in Figure 4-b what this best price is.
a) Performance with one additional experiment.
b)“Best” candidate experiment price $p_{2}$.
Figure 4: Value of one additional global experiment for various $(p_{1},q_{1})$.
Figure 4 shows that by allowing the decision-maker to judiciously select a second price at which to collect information, they can considerably improve the performance of the final price selected. Indeed, while a single price already guarantees a ratio of at least $30\%$ for conversion rates between $[0.01,0.99]$, the decision-maker is ensured to obtain more than $50\%$ of the revenue by experimenting correctly at a second price. More generally, across conversion rates at the initial price, the decision-maker can increase the robust performance by approximately $20\%$ through an additional well-selected price experiment. Figure 4-b also illustrates how our procedure allows us to select a second price to run efficient experimentation. For instance, it provides general guidelines regarding when the price should be deflated or inflated to gather more information: in this case, when the conversion rate at $p_{1}$ is lower (respectively higher) than approximately $0.3$ the decision-maker should consider experimenting at a lower (resp. higher) price.
4.3 Stopping criterion for dynamic pricing with deterministic feedback
In Section 4.2 we considered the case of a decision-maker who explored at one more price before committing, we illustrated how our analysis can provide an initial step towards more efficient dynamic pricing algorithms. In this section, we show how our framework can be used to provide a criterion to stop pricing experiments as soon as possible while guaranteeing a certain level of performance. More precisely, given a desired accuracy level $\epsilon$, a valid stopping criterion ensures that whenever the pricing algorithm is stopped, the pricing mechanism which is outputted will generate at least a $1-\epsilon$ fraction of the optimal revenue. We will illustrate that our framework can be used to provide a stopping criterion on top of any dynamic pricing algorithm used to gather information. To illustrate this, we consider the hypothetical case in which the seller is dynamically adjusting the posted prices and observes deterministic true conversion rates at the offered prices in each round. For instance, we consider the setting in which the value distribution is regular, i.e., $\pazocal{C}=\pazocal{F}$. In this case, given that the revenue function is unimodal, the classical Ternary search algorithm, described in Algorithm 1, can be used for dynamic pricing to ensure one approaches optimal performance.
Input: Acceptable sub-optimality $\epsilon$;
Initialize $a=\underline{v}$, $b=\overline{v}$, $\mathscr{I}_{\mathbf{p},\mathbf{q}}^{1}=\{(\underline{v},1),(\overline{v},0)\}$;
while $t<\frac{\log(\frac{\overline{v}-\underline{v}}{\overline{v}\cdot\epsilon})}{\log(3/2)}$ ;
  // Ternary search usual stopping criterion
 do
       /* Exploration procedure of the Ternary search algorithm */
       Price at $p_{1}=a+\frac{b-a}{3}$, observe $q_{1}$ ;
       Price at $p_{2}=a+\frac{2(b-a)}{3}$, observe $q_{2}$ ;
       Record observed conversion rates $\mathscr{I}_{\mathbf{p},\mathbf{q}}^{t+1}=\mathscr{I}_{\mathbf{p},\mathbf{q}}^{t}\cup\{(p_{1},q_{1}),(p_{2},q_{2})\}$ ;
      if $p_{1}\cdot q_{1}<p_{2}\cdot q_{2}$ then
             Set $a=p_{1}$ ;
       end if
      else
             Set $b=p_{2}$ ;
       end if
      Increment $t$
end while
Return $a$ if $a\cdot q_{a}\geq b\cdot q_{b}$ else return $b$ ;
  // Exploitation price
Algorithm 1 Ternary Search
The ternary search procedure proceeds in steps, sequentially eliminating regions of the price space by leveraging the unimodality of the revenue curve. We note that any dynamic pricing algorithm designed to approach the optimal performance has a structure similar to Algorithm 1. It is defined by a stopping criterion, an exploration procedure and a final exploitation price. In what follows, we fix the exploration procedure of a given dynamic pricing method $\theta$ and combine it with a stopping criterion and an exploitation price designed by leveraging our framework. Recall $\underline{\pazocal{L}}(\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}),\mathbb{A})$ defined in (3.3). Algorithm 2 formalizes our meta-algorithm.
Input: Dynamic pricing algorithm $\theta$; Acceptable sub-optimality $\epsilon$; parameter $M$ for discretization in (3.3);
Initialize $\mathscr{I}_{\mathbf{p},\mathbf{q}}^{1}$;
while $\underline{\pazocal{L}}(\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{t}),\mathbb{A})<1-\epsilon$ do
       Use the exploration procedure of $\theta$;
       Update $\mathscr{I}_{\mathbf{p},\mathbf{q}}^{t+1}$ and increment $t$;
end while
Return $(\psi^{*}_{i})_{i\in\{1,\ldots,M\}}$ solution of the discretized problem (3.3) ;
Algorithm 2 Meta-algorithm for dynamic pricing
Remark that Algorithm 2 provides a valid stropping criterion by design. Indeed, $\underline{\pazocal{L}}(\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{t}),\mathbb{A})$, defined in (3.3), is a lower bound on the maximin ratio. Hence, whenever the algorithm stops, the maximin ratio is at least $1-\epsilon$ and the mechanism returned by (3.3) is ensured to achieve that ratio. We show in Lemma 8 (stated and proved in Section 6) that the usual stopping criterion of the Ternary search algorithm is also valid.
In what follows, we illustrate the effectiveness of our proposed framework in reducing the number of price queries needed to ensure that the worst-case ratio is at least $99\%$. To compare different stopping criteria, we fix the Ternary search algorithm structure, and vary the stopping criterion. In particular, we sample random instances from a parametric family of regular demand curves (either linear demand curves or exponential ones). The linear model satisfies for every $v\in[0,\infty)$,
$\overline{F}(v)=\min\left(\max(a-b\cdot v),0),1\right),$
where $b$ is uniformly sampled from $[1,5]$ and $a$ is uniformly sampled from $[1,b]$. For the exponential demand we have for every $v\in[0,\infty)$,
$\overline{F}(v)=\min(\exp(a-b\cdot v),1),$
where $b$ is uniformly sampled from $[1,5]$ and $a$ is uniformly sampled from $[-0.2,b]$.
For each instance, we report the number of price queries needed to achieve at worst-case ratio greater or equal to $99\%$ when
  1. (i)
    using the stopping criterion of Ternary search which uses $N^{\mathrm{Ternary}}(\epsilon)=2\cdot\left\lceil\frac{\log(\frac{\overline{v}-\underline{v}}{\overline{v}\cdot\epsilon})}{\log(3/2)}\right\rceil$ price experiments, and equal to $44$ in this setting,
  2. (ii)
    using our criterion with $\pazocal{C}=\pazocal{F}$, i.e., when we know that the demand is regular,
  3. (iii)
    using our criterion with $\pazocal{C}=\pazocal{G}$ and adding the constraint that the revenue function is unimodal (this can be implemented by only including constraints in (3.3b) for $i\in\{1,\ldots,M+1\}$ such that $a_{i}\in[a,b]$, where $a$ and $b$ are the variables used by Algorithm 1).
We report in Figure 5 the number of price queries needed by each criterion to ensure that the exploitation price will achieve a $99\%$ worst-case ratio. We emphasize that while the stopping criterion changes, the dynamic pricing algorithm used for exploration is always the same and corresponds to the Ternary search approach.
a) Linear demand.
b) Exponential demand.
Figure 5: Distribution of the number of price queries needed to achieve 99% worst-case ratio performance. ($[\underline{v},\overline{v}]=[1,100]$)
Figure 5 shows that using our characterization of the worst-case ratio allows to stop considerably earlier than by using the standard stopping criterion of Ternary search. When using our stopping criterion which only uses the unimodality of the revenue curve, we are already able to divide by two the number of price queries needed on average. This difference can be explained by the data-driven nature of our criterion. Indeed, the Ternary search stopping criterion is non-adaptive and considers the worst-case number of iterations needed across all unimodal curves, whereas our stopping criterion leverages the information gathered at each step to rule out some of the unimodal curves based on the conversion rates observed.
Furthermore, our stopping criterion can be used to incorporate additional beliefs about the demand distribution. For instance, we can incorporate regularity by setting $\pazocal{C}=\pazocal{F}$. In that case, we see that our criterion can reduce on average the number of price points at which we experiment by a factor of five compared to the Ternary search criterion. For some of the instances considered our regular stopping criterion proposes to stop after only 5 price queries (as opposed to $44$ with usual stopping rules and ternary search). This significant gain in efficiency considerably changes our understanding of the value of dynamic price experiments.
5 Conclusion
We present in this work a framework to precisely quantify the value of data obtained through price experiments in the form of past offered prices and observed conversion rates. Our main methodological contribution is a reduction of an optimization problem whose value quantifies the worst-case revenue gap incurred by a decision-maker who only has access to certain conversion rates compared to an oracle knowing the full value distribution. We leverage this reduction to derive insights on the value of diverse types of historical data but also to prescribe efficient ways of running price experiments.
Throughout this work, we showed that local experiments (such as the sign or the value of the gradient of the revenue function at an offered price) may already encode information useful for pricing purposes. We also demonstrate that our exact quantification of the value of these different types of information provides a tool for rigorous cost-benefit analyses. Indeed, decision-makers may assess whether the benefits associated with obtaining a precise value of the gradient compared to its sign justify the cost of running a longer experiment.
We also illustrate how our framework can be used to improve price experiments. We first propose a procedure to select the next price at which one should experiment before committing to a final price and we quantify the value associated with such decisions. Finally, we show that our framework allows to significantly reduce the length of price experiments even when committing to standard dynamic algorithms (such as Ternary search) by providing a data-driven stopping criterion tailored to the structure of the pricing problem. This stopping criterion is adaptive to the the information collected.
We believe that this work opens up many exciting research avenues towards more efficient price experimentation. One natural question would be to extend our result to be applicable to quantify the value of noisy observations of the demand. Furthermore, the use of our method to prescribe the last price at which to experiment before committing to a final price (as discussed in Section 4.2) provides an interesting first step towards more efficient algorithms for dynamic pricing. We hope that our work inspires new algorithmic principles to trade-off exploration and exploitation in dynamic pricing over multiple steps.
Acknowledgment
The authors would like to express their sincere gratitude to Amine Allouah for his invaluable feedback and insightful discussions during the early stages of this work.
6 Auxiliary Results
Proposition 3 (Feasible information sets).
Let $\mathscr{I}_{\mathbf{p},\mathbf{q}}=\{(p_{i},q_{i}):i=0\ldots,N+1\}$ be a set of sorted increasing prices and associated ccdf values with $(p_{0},q_{0})=(\underline{v},1)$ and $(p_{N+1},q_{N+1})=(\overline{v},0)$. Then:
  • $\pazocal{G}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-empty if and only if the sequence $(q_{i})_{0\leq i\leq N+1}$ is non-increasing.
  • $\pazocal{F}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-empty if and only if the sequence $(q_{i})_{0\leq i\leq N+1}$ is non-increasing and the sequence of slopes $\left(\frac{\Gamma^{-1}\left(q_{i+1}\right)-\Gamma^{-1}\left(q_{i}\right)}{p_{i+1}-p_{i}}\right)_{0\leq i\leq N-1}$ is non-decreasing.
Proof of Proposition 3.
Let $\mathscr{I}_{\mathbf{p},\mathbf{q}}=\{(p_{i},q_{i}):i=0\ldots,N+1\}$ a set of sorted increasing prices and associated ccdf values with $(p_{0},q_{0})=(\underline{v},1)$ and $(p_{N+1},q_{N+1})=(\overline{v},0)$.
Case $\pazocal{C}=\pazocal{F}:$
$\Longrightarrow$) If $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ is $\pazocal{F}$-feasible then there exists a distribution characterized by its cdf $F$ in $\pazocal{F}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Then by definition of a cdf, the sequence $(q_{i}=\overline{F}(p_{i}))_{0\leq i\leq N+1}$ is non-increasing. Fix $11\leq i\leq N-$ 1, Using Lemma 1 on $\overline{F}$ with $(s,q_{s})=(p_{i-1},q_{i-1})$ and $(s^{\prime},q_{s^{\prime}})=(p_{i+1},q_{i+1})$, we have
$\overline{F}(p_{i})=q_{i}\geq\Gamma\left(\Gamma^{-1}\left(q_{i-1}\right)+\frac{\Gamma^{-1}\left(q_{i+1}\right)-\Gamma^{-1}\left(q_{i-1}\right)}{p_{i+1}-p_{i-1}}(p_{i}-p_{i-1})\right)$
which implies that $\frac{{\Gamma^{-1}\left(q_{i}\right)}-{\Gamma^{-1}\left(q_{i-1}\right)}}{{p_{i}}-{p_{i-1}}}\leq\frac{{\Gamma^{-1}\left(q_{i+1}\right)}-{\Gamma^{-1}\left(q_{i-1}\right)}}{{p_{i+1}}-{p_{i-1}}}.$ Furthermore, note that
$\frac{{\Gamma^{-1}\left(q_{i+1}\right)}-{\Gamma^{-1}\left(q_{i-1}\right)}}{{p_{i+1}}-{p_{i-1}}}=\frac{\Gamma^{-1}(\overline{F}(p_{i+1}))-\Gamma^{-1}(\overline{F}(p_{i-1}))}{{p_{i+1}}-{p_{i-1}}}\leq\frac{\Gamma^{-1}(\overline{F}(p_{i+1}))-\Gamma^{-1}(F(p_{i}))}{{p_{i+1}}-p_{i}}=\frac{{\Gamma^{-1}\left(q_{i+1}\right)}-{\Gamma^{-1}\left(q_{i}\right)}}{{p_{i+1}}-{p_{i}}},$
where the inequality holds because $\Gamma^{-1}(\overline{F}(\cdot))$ is convex (see proof of Lemma 2 in Allouah et al. (2022)) and the slope function of a convex function is non-decreasing.
Therefore, $\frac{{\Gamma^{-1}\left(q_{i}\right)}-{\Gamma^{-1}\left(q_{i-1}\right)}}{{p_{i}}-{p_{i-1}}}\leq\frac{{\Gamma^{-1}\left(q_{i+1}\right)}-{\Gamma^{-1}\left(q_{i}\right)}}{{p_{i+1}}-{p_{i}}}$, which proves the implication.
[…]
Case $\pazocal{C}=\pazocal{G}:$
$\Longrightarrow$) If $\mathscr{I}_{\mathbf{p},\mathbf{q}}$ is $\pazocal{G}$-feasible then there exists a distribution characterized by its cdf $F$ in $\pazocal{G}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Then by definition of a cdf, the sequence $(q_{i}=\overline{F}(p_{i}))_{0\leq i\leq N+1}$ is non-increasing.
[…]
Lemma 6.
Fix three scalars $\alpha,\beta>0$ and $s\geq 0$ and a cumulative distribution function $F$ such that $\overline{F}=\Gamma\left(\alpha+\beta(v-s)\right)$ for $v\geq s$. Then $F$ admits a constant virtual value function equal to $s-\frac{1+\alpha}{\beta}.$
Proof of Lemma 6.
Let us first explicitly compute the virtual value function. The derivative of $\Gamma\left(\alpha+\beta(v-s)\right)$, for any $v\geq s$, is equal to $-\beta\Gamma\left(\alpha+\beta(v-s)\right)^{2}.$ Therefore, the virtual value function evaluated at $v\geq s$ is equal to,
$v-\frac{\Gamma\left(\alpha+\beta(v-s)\right)}{\beta\Gamma\left(\alpha+\beta(v-s)\right)^{2}}=v-\frac{1}{\beta\Gamma\left(\alpha+\beta(v-s)\right)}=v-\frac{1+\alpha+\beta(v-s)}{\beta}=s-\frac{1+\alpha}{\beta}.\qed$
Lemma 7.
For any $r^{*}$ such that ${\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ is non-empty, $\sup{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})=\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})$.
Proof of Lemma 7.
Step 1: We first show that, $\sup{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\leq\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})$. Let $q\in{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ and let $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ such that ${r^{*}}$ is an optimal price for $F$ with associated conversion rate $q$. Consider $i\in\{0,\ldots,N\}$ such that ${r^{*}}\in[p_{i},p_{i+1})$ The crossing lemma implies that, $q=\overline{F}({r^{*}}-)\leq\min\left(\overline{G}_{\pazocal{F},{i-1}}({r^{*}}-),\overline{G}_{\pazocal{F},{i+1}}({r^{*}}-)\right)\leq\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}}),$ where the last inequality holds by definition of $\overline{U}_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ when $\pazocal{C}$ is either $\pazocal{G}$ or $\pazocal{F}$.
Step 2: Next, we show that $\sup{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})\geq\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})$ by proving that $\overline{U}_{\pazocal{C}}({r^{*}}|\mathscr{I}_{\mathbf{p},\mathbf{q}})\in{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$. To do so, we need to show that there exists a distribution $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ such that ${r^{*}}$ is an optimal price and $\overline{F}({r^{*}}-)=\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})$. We consider as a candidate the distribution $F_{\pazocal{C}}(\cdot|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}})$ defined in Equation 2.8. Let $\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+}=\mathscr{I}_{\mathbf{p},\mathbf{q}}\cup({r^{*}},\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}}))$ and note that, $U_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}})\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$ and therefore this set is non-empty. Hence, by Lemma 3, $L_{\pazocal{C}}(\cdot|\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})$, which implies that, $F_{\pazocal{C}}(\cdot|{r^{*}},\mathscr{I}_{\mathbf{p},\mathbf{q}})\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}}^{+})\subset\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$.
To show that ${r^{*}}$ is the optimal price, we see from Lemma 4 that it suffices to show that ${r^{*}}\cdot\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})\geq\max_{i}p_{i}\cdot q_{i}$. Let $q\in{\mathscr{}B}_{r^{*}}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ then by definition there exists $F\in\pazocal{C}(\mathscr{I}_{\mathbf{p},\mathbf{q}})$ such that $\overline{F}({r^{*}}-)=q$ and ${r^{*}}\cdot\overline{F}({r^{*}}-)\geq p\cdot\overline{F}(p-)$ for every $p\in[\underline{v},\overline{v}]$. This implies that for every $i\in\{0,N+1\}$,
${r^{*}}\cdot\overline{U}_{\pazocal{C}}({r^{*}}-|\mathscr{I}_{\mathbf{p},\mathbf{q}})\stackrel{{\scriptstyle(a)}}{{\geq}}{r^{*}}\cdot q={r^{*}}\cdot\overline{F}({r^{*}}-)\geq p_{i}\cdot\overline{F}(p_{i}-)=p_{i}\cdot q_{i},$
where $(a)$ follows from step 1. This concludes the proof. ∎
Lemma 8.
The ternary stopping criterion is a valid stopping criterion.
Proof of Lemma 8..
Denote by $T_{N}=[a_{N},b_{N}]$, the interval obtained at iteration $N$ in Algorithm 1, by uni-modality of $\mathrm{Rev}\left(\cdot|F\right)$ we know that $r_{F}$ is in $[a_{N},b_{N}]$, therefore we have:
$1-\frac{\mathrm{Rev}\left(a_{N}|F\right)}{\mathrm{Rev}\left(r_{F}|F\right)}$ $=\frac{r_{F}\cdot\overline{F}(r_{F}-)-a_{N}\cdot\overline{F}(a_{N}-)}{\mathrm{Rev}\left(r_{F}|F\right)}$
$\leq\frac{(r_{F}-a_{N})\cdot\overline{F}(r_{F}-)}{\mathrm{Rev}\left(r_{F}|F\right)}\leq\frac{b_{N}-a_{N}}{\mathrm{Rev}\left(r_{F}|F\right)}=\frac{\overline{v}-\underline{v}}{\mathrm{Rev}\left(r_{F}|F\right)}\left(\frac{2}{3}\right)^{N}\leq\frac{\overline{v}-\underline{v}}{\underline{v}}\left(\frac{2}{3}\right)^{N}$
Therefore for $N\geq N(\epsilon)=\frac{\log(\frac{\overline{v}-\underline{v}}{\underline{v}\cdot\epsilon})}{\log(3/2)}$, we obtain the desired level of the maximin ratio. ∎